### v2

# Curvature pinching of asymptotically conical gradient expanding Ricci solitons

Abstract.
Since the well-known work of Hamilton [62] and Ivey [64], the Hamilton-Ivey curvature pinching and its generalizations have become a signature feature of gradient shrinking and steady Ricci solitons, and more generally, of ancient solutions to the Ricci flow. However, analogous results for gradient expanding Ricci solitons have remained elusive. This is largely due to the fact that the proofs of existing curvature pinching estimates crucially rely on shrinking and steady solitons being ancient, a property not shared by gradient Ricci expanders. In this paper, we investigate curvature pinching phenomena in non-compact asymptotically conical gradient expanding Ricci solitons and establish several Hamilton-Ivey type curvature pinching estimates. These results are parallel to those known for shrinking and steady Ricci solitons. In particular, we prove a three-dimensional Hamilton-Ivey type curvature pinching theorem: any three-dimensional non-compact gradient Ricci expander, which is asymptotic to a cone with positive scalar curvature, must have positive sectional curvature. As an application, we combine our result with that of Deruelle [51] to establish a uniqueness theorem for three-dimensional asymptotically conical expanders with positive scalar curvature.
Furthermore, we formulate a general proof method and apply it to obtain analogues of several additional known generalized Hamilton-Ivey type curvature pinching results for ancient solutions. Among these is a curvature pinching estimate for four-dimensional asymptotically conical Ricci expanders with uniformly positive isotropic curvature, analogous to a result for four-dimensional gradient steady solitons due to Brendle [7].
1. Introduction
A complete Riemannian manifold $(M^{n},g)$ is said to be a gradient Ricci soliton if there exists a smooth function $f$ on $M^{n}$ such that the Ricci tensor $Rc$ of the metric $g$ satisfies the equation
$Rc+\nabla^{2}f=\rho g$ (1.1)
for some constant $\rho\in{\mathbb{R}}$, where $\nabla^{2}f$ denotes the Hessian of $f$. The Ricci soliton is said to be expanding, or steady, or shrinking if $\rho<0$, or $\rho=0$, or $\rho>0$. The function $f$ is called a potential function of the gradient Ricci soliton.
Gradient Ricci solitons generate self-similar solutions to Hamilton’s Ricci flow [59] and play an important role in the study of the formation of singularities [62, 79]. In particular, shrinking and steady solitons often arise as Type I and, respectively, Type II singularity models [62, 76, 55, 16] in the Ricci flow, while expanding solitons may arise as Type III singularity models [16, 36] and over which the matrix Li-Yau-Hamilton (LYH) differential Harnack inequality [61, 14] becomes equality.
The first examples of gradient expanding Ricci solitons are the one-parameter family of complete, rotationally symmetric, asymptotically conical gradient expanding Ricci solitons on $\mathbb{R}^{n}$ ($n\geq 3$) with positive or negative sectional curvature constructed by Bryant [11], and the one-parameter family of complete, $U(n)$-invariant, asymptotically conical gradient expanding Kähler-Ricci solitons on $\mathbb{C}^{n}$ with similar curvature behavior constructed by the first author [16]. The constructions in [15, 16] were later extended by Feldman-Ilmanen-Knopf [56] to produce gradient expanding Kähler-Ricci solitons on the complex line bundles O(k)O(-k) ($k>n$) over the complex projective space ${\mathbb{C}}P^{n}$ ($n\geq 1$), and further generalized by Dancer-Wang [50]. Additional constructions can be found in [65, 2, 49, 57, 12, 85, 86, 77, 41, 42].
Asymptotically conical gradient expanding Ricci solitons have received increasing attention in recent years. Chodosh [44] showed that any gradient expanding Ricci soliton with positive sectional curvature that is asymptotic to a Euclidean cone must be rotationally symmetric. A similar result for gradient Kähler-Ricci expanders was obtained by Chodosh-Fong [45]. Schulze-Simon [82] constructed gradient expanding solitons emerging from the asymptotic cones at infinity of Ricci flow solutions on complete, non-compact, Riemannian manifolds with bounded, nonnegative curvature operator and positive asymptotic volume ratio. Deruelle [51] proved that any Riemannian cone whose link is a differentiable sphere with curvature operator $Rm>1$ can be smoothed out by the Ricci flow into a gradient expanding Ricci soliton with nonnegative curvature operator. In the Kähler setting, Conlon, Deruelle, and Sun [47, 48] established the existence and uniqueness of asymptotically conical gradient expanding Kähler-Ricci solitons on smooth canonical models of Kähler cones.
More recently, Chan-Lee-Peachey [33] showed that any metric cone at infinity of a non-collapsed weakly PIC1 manifold is resolved by a gradient expanding Ricci soliton. Bamler-Chen [4] developed a degree theory for 4-dimensional, asymptotically conical expanders, which implies the existence of gradient expanders asymptotic to any cone over $\mathbb{S}^{3}$ with nonnegative scalar curvature. Additionally, Chan-Lee [32] constructed various examples of asymptotically conical gradient expanders with positive curvature and exotic curvature decay. In particular, [58, 1, 4] have proposed that asymptotically conical gradient Ricci expanders may serve to continue Ricci flow past singular time and resolve conical singularities. For other related developments, see [46, 28, 81, 5, 37, 29, 83, 53, 52, 69] and the references therein.
Curvature estimates for gradient expanding Ricci solitons with $Rc\geq 0$ or scalar curvature $R>0$ have also been established in [22, 23], mirroring those for shrinking solitons [74, 75] or steady solitons [21, 30, 17].
Despite the progress described above, a key feature well known for gradient shrinking and steady Ricci solitons has remained absent in the expanding case: Hamilton-Ivey type curvature pinching. As is well-known, a hallmark of the three-dimensional Ricci flow is the Hamilton-Ivey curvature pinching [62, 64] (see also [27, Theorem 2.4.1]), which asserts that when curvature blows up, the positive part blows up at a faster rate than the absolute value of the negative part. In particular, 3-dimensional singularity models that are shrinking or steady gradient Ricci solitons, or more generally ancient solutions, must have nonnegative sectional curvature. This remarkable feature is powerful: it enables the application of the Li-Yau-Hamilton differential Harnack inequality and special geometry of nonnegatively curved 3-manifolds to effectively analyze three-dimensional singularity models. The Hamilton-Ivey curvature pinching was later extended by B.-L. Chen [35] to arbitrary 3-dimensional ancient solutions [^1] Ancient solutions exist for all past time, up to a final singular time; immortal solutions, starting at some initial time, exist for all future time., which include gradient shrinking and steady solitons as important special cases. It implies that any complete ancient solution in dimension three must have nonnegative sectional curvature. This has played a crucial role in the classifications of 3-dimensional gradient shrinking and steady Ricci solitons [20, 6, 66, 67, 68], as well as 3D ancient solutions [8, 9].
In higher dimensions, various forms of the generalized Hamilton-Ivey curvature pinching have been established for ancient solutions under suitable assumptions, see, e.g., [89, 3, 70, 43]. More recently, the authors proved certain curvature pinching properties for $4$-dimensional ancient solutions with positive isotropic curvature (PIC) or half PIC [24, 25], leading to partial classifications of $4$-dimensional shrinking and steady gradient Ricci solitons with weakly positive isotropic curvature (WPIC) or half WPIC.
However, analogous curvature pinching results for gradient expanding Ricci solitons have remained largely unknown. This is partly because the proofs of Hamilton-Ivey-type estimates for shrinking and steady solitons rely in an essential way on the ancient nature of these solutions. Since gradient expanders are not ancient solutions but special immortal solutions, those methods do not apply.
Motivated in part by the second author’s work [87] on the convexity of mean convex, asymptotically conical self-expanders in mean curvature flow, and in part by the close resemblance between curvature estimates for Ricci expanders with nonnegative curvature [22, 23] and those for shrinking or steady Ricci solitons [74, 75, 21, 30, 17] (see especially the comparison in dimension four given in [17]), we began to investigate whether analogous Hamilton-Ivey-type curvature pinching properties might hold in the expanding case.
By adapting an argument from the second author’s recent work [87] on asymptotically conical mean curvature expanders – inspired in turn by Spruck-Xiao [84] and Xie-Yu [88] – we have indeed established a number of generalized Hamilton-Ivey-type curvature pinching results for non-compact, asymptotically conical, gradient expanding Ricci solitons with positive scalar curvature.
Our first result is a Hamilton-Ivey type curvature pinching for $3$-dimensional non-compact, asymptotically conical, gradient expanders with positive scalar curvature.
Theorem 1.1.
Let $(M^{3},g,f)$ be a $3$-dimensional non-compact, asymptotically conical gradient expanding Ricci soliton. Suppose the asymptotic cone has positive scalar curvature. Then, $(M^{3},g,f)$ must have positive sectional curvature.
Remark 1.1.
Theorem 1.1 may be viewed as an analogue of B.-L. Chen’s result for three-dimensional ancient solutions to the Ricci flow [35]. Moreover, after our paper was completed, the authors learned from P.-Y. Chan that a similar result to Theorem 1.1 was proved in [34, Corollary 1.16] by a different method.
As an application of Theorem 1.1, for $n=3$, we can extend Deruelle’s uniqueness result [51, Theorem 1.3] to asymptotically conical expanding Ricci solitons with positive scalar curvature.
Corollary 1.1.
Let $(\Sigma^{2},\bar{g})$ be a smooth 2-sphere with Gaussian curvature $\bar{K}>1$. Then, there exists a unique $3$-dimensional gradient expanding Ricci soliton with positive scalar curvature asymptotic to the cone $\mathcal{C}^{3}:=[0,\infty)\times\Sigma^{2}$ with link $\Sigma^{2}$.
For $n\geq 4$, we have the following Hamilton-Ivey type curvature pinching result for asymptotically conical, gradient expanding solitons with positive scalar curvature and vanishing Weyl tensor.
Theorem 1.2.
Let $(M^{n},g,f)$, $n\geq 4$, be an $n$-dimensional non-compact, locally conformally flat, asymptotically conical gradient expanding Ricci soliton. Suppose the asymptotic cone has positive scalar curvature. Then, $(M^{n},g,f)$ must have positive curvature operator.
Remark 1.2.
Theorem 1.2 is an analogue of Z.-H. Zhang’s result for locally conformally flat gradient shrinking (and steady) solitons [89]. Moreover, by the work of [19, 26], the locally conformally flat assumption can be replaced by the weaker assumption of vanishing $D$-tensor introduced in [18, 19].
As a consequence, by combining with the work of Cao-Chen [19] and Chen-Wang [39], we have the following application in dimension four.
Corollary 1.2.
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact, half conformally flat, asymptotically conical gradient expanding Ricci soliton. Suppose the asymptotic cone has positive scalar curvature. Then, $(M^{4},g,f)$ has positive curvature operator.
Our next two results concern curvature pinching of 4-dimensional asymptotically conical gradient expanding Ricci solitons with either positive isotropic curvature (PIC) or half positive isotropic curvature (half-PIC); see Section 2.2 for definitions.
Theorem 1.3.
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with half positive isotropic curvature. If the asymptotic cone has positive scalar curvature and satisfies either $A\geq 0$ or $C\geq 0$, then $(M^{4},g,f)$ has either $A>0$ or $C>0$.
Theorem 1.4.
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with positive isotropic curvature. If the asymptotic cone has positive scalar curvature, then the Ricci curvature of $(M^{4},g,f)$ is $2$-positive.
Remark 1.3.
Theorems 1.3 and 1.4 form an analogue of the authors’ previous results for $4$-dimensional complete ancient solutions with half PIC and PIC [24, 25], respectively. Moreover, both results are valid under some slightly weaker assumptions; see Theorem 4.1 and Theorem 4.2 in Section 4.
By observing common patterns in the proofs of Theorem 1.3 and Theorem 1.4, we formulate a general method of proof (Lemma 5.1) and apply it to obtain analogues of several additional known generalized Hamilton-Ivey type curvature pinching results for ancient solutions. This includes the following result for $4$-dimensional asymptotically conical gradient expanding Ricci soliton with uniformly positive isotropic curvature (UPIC), as well as several others stated in Theorem 5.1.
Theorem 1.5.
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with uniformly positive isotropic curvature. If the asymptotic cone is a non-flat Euclidean cone, then the curvature operator of $(M^{4},g,f)$ is positive.
Remark 1.4.
Theorem 1.5 is an analogue of Brendle’s curvature pinching result for $4$-dimensional gradient steady Ricci solitons with UPIC [7], as well as of Cho-Li’s result for $4$-dimensional complete ancient solutions with UPIC [43].
By combining Theorem 1.5 with Chodosh’s work [44], we obtain the following classification for 4D asymptotically conical expanding Ricci solitons with UPIC.
Corollary 1.3.
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact gradient expanding Ricci soliton with uniformly positive isotropic curvature. If $(M^{4},g,f)$ is asymptotic to a non-flat Euclidean cone, then it is rotationally symmetric.
Organization of the Paper. Section 2 introduces the notation and basic concepts used throughout the paper, and collects several useful facts needed for the main arguments. Section 3 is devoted to the proofs of Theorem 1.1 and Theorem 1.2. In Section 4, we present the proofs of Theorem 1.3 and Theorem 1.4. In Section 5, we formulate a general lemma that can be applied especially to asymptotically conical gradient expanding Ricci solitons and use it to prove Theorem 1.5 and Theorem 5.1. Finally, the Appendix contains some elementary curvature properties of cones used in Section 3 and Section 4.
Acknowledgements. The first author was partially supported by a Simons Fellowship and a grant from the Simons Foundation. The second author would like to thank Prof. Xiaochun Rong for his continual support and encouragement. Both authors would like to thank Dr. P.-Y. Chan for very helpful information that allowed the removal of the positive scalar curvature assumption on $(M^{n},f,g)$ in Theorem 1.1 and Theorem 1.2 in an earlier version of the paper.
2. Preliminaries
In this section, we fix notation and recall several basic facts and results that will be used throughout the paper. Throughout, we denote by
$Rm=\{R_{ijkl}\},\quad Rc=\{R_{ij}\},\quad R$
the Riemann curvature tensor, the Ricci tensor, and the scalar curvature of the metric $g=\{g_{ij}\}$, respectively, either in local coordinates or with respect to a local orthonormal frame.
2.1. Asymptotically conical expanding Ricci solitons
Recall that, in general, by an $n$-dimensional cone we mean an $n$-manifold
$\mathcal{C}:=[0,\infty)\times\Sigma^{n-1}$
equipped with the Riemannian metric
$g_{c}=dr^{2}+r^{2}\bar{g}_{\Sigma},$
where $(\Sigma^{n-1},\bar{g}_{\Sigma})$, called the link of the cone $\mathcal{C}$, is a closed (n1)(n-1)-dimensional Riemannian manifold. As an example, the standard non-flat Euclidean cone with cone angle $\alpha\in[0,1)$ is given by the conical metric $g_{\alpha}$ on $\mathbb{R}^{n}\setminus\{0\}$, expressed in polar coordinates as
$g_{\alpha}:=dr^{2}+(1-\alpha)r^{2}\bar{g}_{\mathbb{S}^{n-1}(1)}.$
In general, for any cone $\mathcal{C}$ over the link $\Sigma$ and for $s\geq 0$, set
$E_{s}=(s,\infty)\times\Sigma\subset\mathcal{C},$
and define the dilation by $\tau>0$ as the map
$\rho_{\tau}:E_{0}\to E_{0},\qquad\rho_{\tau}(r,\sigma)=(\tau r,\sigma).$
Definition 2.1.
  • (a)
    A Riemannian manifold $(M^{n},g)$ is said to be $C^{k}$-asymptotic to the cone $(E_{0},g_{c})$ if, for some $s>0$, there exists a diffeomorphism
    $\Phi:E_{s}\to M\setminus K,$
    for some compact subset $K\subset M$, such that
    $\tau^{-2}\rho_{\tau}^{*}\Phi^{*}g\longrightarrow g_{c}\quad\text{as }\tau\to\infty\quad\text{in }C^{k}_{\mathrm{loc}}(E_{0},g_{c}).$
  • (b)
    We say that a Riemannian manifold $(M,g)$ is asymptotically conical if there exists a cone $(E_{0},g_{c})$ such that $(M,g)$ is $C^{k}$-asymptotic to $(E_{0},g_{c})$ for all integers $k\geq 0$.
2.2. Curvature decomposition and isotropic curvature of four-manifolds
In this subsection, we recall some facts about the curvature decomposition and isotropic curvature of $4$-manifolds. For more background, we refer the reader to Hamilton’s paper [63] and our previous work [24, 25].
For any oriented Riemannian $4$-manifold $(M^{4},g)$, the bundle of $2$-forms admits the decomposition
$\wedge^{2}(M)=\wedge^{+}(M)\oplus\wedge^{-}(M),$
where $\wedge^{+}(M)$ and $\wedge^{-}(M)$ denote the subbundles of self-dual and anti-self-dual $2$-forms, respectively. With respect to this splitting, the curvature operator has block form
${Rm}=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\tfrac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\tfrac{R}{12}I\end{pmatrix}$,
[…]
Let $A_{1}\leq A_{2}\leq A_{3}$ and $C_{1}\leq C_{2}\leq C_{3}$ denote the eigenvalues of $A$ and $C$, respectively, and let $a_{1}\leq a_{2}\leq a_{3}$ and $c_{1}\leq c_{2}\leq c_{3}$ denote the eigenvalues of $W^{+}$ and $W^{-}$, respectively. Then
$A_{i}=a_{i}+\tfrac{R}{12},\qquad C_{i}=c_{i}+\tfrac{R}{12},\qquad(i=1,2,3).$
Since $a_{1}+a_{2}+a_{3}=\operatorname{tr}W^{+}=0=\operatorname{tr}W^{-}=c_{1}+c_{2}+c_{3}$, it follows that
$\operatorname{tr}A=\operatorname{tr}C=\tfrac{R}{4}.$
Definition 2.2.
An $n$-dimensional Riemannian manifold $(M^{n},g)$, $n\geq 4$, is said to have positive isotropic curvature (PIC) if
$R_{1313}+R_{1414}+R_{2323}+R_{2424}-2R_{1234}>0$
for every orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}$. Similarly, it has nonnegative isotropic curvature, or weakly PIC (WPIC) if, for every such frame,
$R_{1313}+R_{1414}+R_{2323}+R_{2424}-2R_{1234}\geq 0.$
The notion of isotropic curvature was first introduced by Micallef-Moore [73], in which they proved that any compact simply connected $n$-dimensional Riemannian manifold with PIC is homeomorphic to a round sphere. It also plays a key role in the convergence theory for the higher dimensional Ricci flow, especially in Brendle-Schoen’s proof of the 1/4-pinching differentiable sphere theorem [10].
It turns out that, in dimension four, these curvature conditions (and their natural extensions, half PIC or half WPIC) can be characterized in terms of the $3\times 3$ matrices $A$ and $C$ as follows:
  • PIC (WPIC) if and only if $A$ and $C$ are 2-positive (weakly 2-positive), i.e., $A_{1}+A_{2}>0$ ($A_{1}+A_{2}\geq 0$) and $C_{1}+C_{2}>0$ ($C_{1}+C_{2}\geq 0$) on $M$ [61];
  • half PIC (half WPIC) if and only if either $A$ or $C$ is 2-positive (weakly 2-positive), i.e., $A_{1}+A_{2}>0$ ($A_{1}+A_{2}\geq 0$) or $C_{1}+C_{2}>0$ ($C_{1}+C_{2}\geq 0$).
Definition 2.3.
$(M^{4},g)$ is said to be uniformly PIC (UPIC) if $M^{4}$ has PIC and in addition satisfies the pointwise pinching condition
$\max\{A_{3},B_{3},C_{3}\}\leq\Lambda\min\{A_{1}+A_{2},C_{1}+C_{2}\}$
on $M^{4}$ for some constant $\Lambda\geq 1$.
2.3. Basic differential equations satisfied by curvatures of Ricci solitons
As a special case of curvature evolution equations under the Ricci flow [60, 63], we have the following well-known curvature differential equations for any gradient Ricci soliton satisfying (1.1).
Lemma 2.1 (cf. Hamilton [60]).
Let $(M^{4},g(t))$ be a $4$-dimensional complete gradient Ricci soliton satisfying Eq. (1.1). Then,
$\Delta_{f}R=2\rho R-2|Rc|^{2},$
$\Delta_{f}Rm=2\rho Rm-2(Rm^{2}+Rm^{\sharp}),$
$\Delta_{f}A=2\rho A-2(A^{2}+2A^{\sharp}+B{}^{t}B),$
$\Delta_{f}B=2\rho B-2(AB+BC+2B^{\sharp}),$
$\Delta_{f}C=2\rho C-2(C^{2}+2C^{\sharp}+{}^{t}BB).$
Here, for any $3\times 3$ matrix $D$, we denote by $D^{2}$ its square and by $D^{\sharp}$ the transpose of its adjoint. In addition, $\Delta_{f}:=\Delta-\nabla f\cdot\nabla$ denotes the weighted Laplace operator.
Remark 2.1.
Except for the first identity in Lemma 2.1, the factor 22 differs from [60] due to our normalization of the inner product on $\wedge^{2}(M)$ (see [24, (2.4)]). Moreover, the first two equations are valid in all dimensions.
2.4. Calabi’s barrier maximum principle
Finally, we shall need the following barrier maximum principle due to Calabi.
Lemma 2.2 ([13]).
Let $\Omega\subset M$ be a bounded connected domain with smooth boundary, and let $u\in C^{0}(\Omega)$. Let $L$ be a uniformly elliptic operator with continuous coefficients and vanishing constant term. If $L(u)\leq 0$ (resp. $L(u)\geq 0$) on $\Omega$ in the barrier sense, then
$\sup_{\Omega}u\geq\sup_{\partial\Omega}u\qquad\bigl(\text{resp. }\inf_{\Omega}u\leq\inf_{\partial\Omega}u\bigr).$
Moreover, if $u$ attains an interior minimum (resp. maximum), then $u$ is a constant in $\Omega$.
3. Curvature pinching for asymptotically conical Ricci expanders with vanishing Weyl tensor
This section is devoted to the proof of Theorem 1.1 and Theorem 1.2. For the reader’s convenience, we restate Theorem 1.2 in the following form.
Theorem 3.1.
Let $(M^{n},g,f)$, $n\geq 4$, be an $n$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with vanishing Weyl tensor $W\equiv 0$. Suppose the asymptotic cone has positive scalar curvature. Then, $(M^{n},g,f)$ has positive curvature operator.
First of all, since the asymptotic cone of $(M^{n},g,f)$ has positive scalar curvature, it follows from the strong maximum principle and [31, Theorem 1.6] that $(M^{n},g,f)$ itself has positive scalar curvature $R>0$. This fact will be used in the proofs of Theorem 3.1 and Theorem 1.1.
Next, shall need the following differential inequality for the ratio between the smallest Ricci eigenvalue and the scalar curvature.
Lemma 3.1 ([54, 80]).
Let $(M^{n},g,f)$, $n\geq 3$, be an $n$-dimensional gradient Ricci soliton with vanishing Weyl tensor $W\equiv 0$ and positive scalar curvature $R>0$. Then, in the barrier sense, we have
$\Delta_{F}\left(\frac{\lambda_{1}}{R}\right)\leq\frac{2h_{1}}{(n-1)(n-2)R^{2}}\leq 0,\qquad(F=f-2\log R)$
where
$\begin{split}h_{1}&=(n-2)\lambda_{1}^{2}(n\lambda_{1}-R)+\bigl((n-2)\lambda_{1}-R\bigr)\Biggl((n-1)\sum_{j=2}^{n}\lambda_{j}^{2}-\Bigl(\sum_{j=2}^{n}\lambda_{j}\Bigr)^{2}\Biggr).\end{split}$
For the reader’s convenience, we also include a proof here.
Proof.
Fix any point $p_{0}\in M$, and choose a unit vector field $e_{1}$ such that $Rc(e_{1})=\lambda_{1}e_{1}$ at $p_{0}$, where $\lambda_{1}$ denotes the smallest Ricci eigenvalue. Extend $e_{1}$ by parallel transport along geodesics emanating from $p_{0}$. Clearly, $\lambda_{1}\leq Rc(e_{1},e_{1})$ with equality at $p_{0}$. Evaluating at $p_{0}$, we obtain, in the barrier sense (cf. Calabi [13]),
$\begin{split}\Delta_{f}\lambda_{1}&\leq\Delta_{f}Rc(e_{1},e_{1})\\ &=(\Delta_{f}Rc)(e_{1},e_{1})\\ &=2\rho\lambda_{1}-\frac{2nR}{(n-1)(n-2)}\lambda_{1}+\frac{4}{n-2}\lambda_{1}^{2}-\frac{2}{n-2}\left(|Rc|^{2}-\frac{R^{2}}{n-1}\right),\end{split}$ (3.1)
where in the last equality, we have used [54, (2.8)] (see also [80, Lemma 2.5]).
On the other hand, we have
$\Delta_{f}\left(\frac{\lambda_{1}}{R}\right)=\frac{1}{R}\Delta_{f}\lambda_{1}-\frac{\lambda_{1}}{R^{2}}\Delta_{f}R-\frac{2}{R^{2}}\langle\nabla\lambda_{1},\nabla R\rangle+\frac{2\lambda_{1}}{R^{3}}|\nabla R|^{2}.$ (3.2)
Let $F=f-2\log R$. Substituting (3.1) and the formula for $\Delta_{f}R$ in Lemma 2.1 into (3.2), we obtain
$\Delta_{F}\left(\frac{\lambda_{1}}{R}\right)\leq\frac{2h_{1}}{(n-1)(n-2)R^{2}},$
where
$h_{1}=R^{2}(R-n\lambda_{1})+(n-1)\Big(2\lambda_{1}^{2}R+\big((n-2)\lambda_{1}-R\big)|Rc|^{2}\Big).$ (3.3)
Now set
$S=\sum_{j=2}^{n}\lambda_{j},\qquad T=\sum_{j=2}^{n}\lambda_{j}^{2}$
so that
$R=\lambda_{1}+S,\qquad|Rc|^{2}=\lambda_{1}^{2}+T.$
Substituting these expressions into (3.3), expanding and regrouping, we obtain
$\begin{split}h_{1}&=S^{2}\big[S+(3-n)\lambda_{1}\big]+(2-n)\lambda_{1}^{2}\big[S-(n-1)\lambda_{1}\big]\\ &\quad+(n-1)T\big((n-3)\lambda_{1}-S\big)\\ &=(n-2)\lambda_{1}^{2}(n\lambda_{1}-R)+\big[(n-2)\lambda_{1}-R\big]\big[(n-1)T-S^{2}\big].\end{split}$
Finally, it is clear from the above expression that $h\leq 0$. ∎
Now, we divide the proof of Theorem 3.1 into two parts.
Part I: $(M^{n},g,f)$ has nonnegativity Ricci curvature $Rc\geq 0$.
Proof of $Rc\geq 0$.
As in [84, 88, 87], we argue by contradiction. Suppose $Rc\ngeq 0$. Then the set
$M^{-}:=\{p\in M:\lambda_{1}(p)<0\}$
is nonempty, hence
$\epsilon_{1}:=\inf_{M}\left(\frac{\lambda_{1}}{R}\right)<0.$
Case 1: Interior infimum. Suppose the negative infimum $\epsilon_{1}$ is attained at some point $p_{0}\in M$. By Lemma 3.1 and Calabi’s barrier strong maximum principle (Lemma 2.2), the ratio $\lambda_{1}/R$ must be constant on some bounded connected domain $p_{0}\in\Omega\subset M^{-}$. In particular, $h\equiv 0$ on $\Omega$ which forces
$\lambda_{1}^{2}(n\lambda_{1}-R)\equiv 0,\qquad\bigl((n-2)\lambda_{1}-R\bigr)\Biggl((n-1)\sum_{j=2}^{n}\lambda_{j}^{2}-\Bigl(\sum_{j=2}^{n}\lambda_{j}\Bigr)^{2}\Biggr)\equiv 0.$ (3.4)
Since $\lambda_{1}<0$ at $p_{0}$ by assumption, from the first equation in (3.4), we must have $R=n\lambda_{1}<0$ at $p_{0}$, contradicting the fact that $R>0$ on $M$. Hence, Case 1 is ruled out.
Case 2: Infimum at infinity. Assume instead that $\lambda_{1}/R$ attains its negative infimum at infinity. Following the argument of [87], there exists a sequence $\{p_{i}\}\subset M$ with $p_{i}\to\infty$ such that
$\lim_{i\to\infty}\frac{\lambda_{1}}{R}(p_{i})=\epsilon_{1}<0.$ (3.5)
Since the asymptotic cone $\mathcal{C}$ of the expanding soliton has positive scalar curvature, the ratio $\tilde{\lambda}_{1}/\tilde{R}$ is well defined on $\mathcal{C}$, where $\tilde{\lambda}_{1}$ and $\tilde{R}$ denote its smallest Ricci eigenvalue and scalar curvature, respectively. For the sequence $\{p_{i}\}\subset M$, one can find a corresponding sequence $\{\tilde{p}_{i}\}\subset\mathcal{C}$ such that
$\lim_{i\to\infty}\frac{\lambda_{1}}{R}(p_{i})=\lim_{i\to\infty}\frac{\tilde{\lambda}_{1}}{\tilde{R}}(\tilde{p}_{i}).$
On the other hand, since the expanding Ricci soliton satisfies $W\equiv 0$, its asymptotic cone $\mathcal{C}$ also has vanishing Weyl tensor. By Lemma 6.2, $\mathcal{C}$ has nonnegative Ricci curvature, whose smallest eigenvalue $\tilde{\lambda}_{1}\equiv 0$ occurs in the radial direction. Hence
$\lim_{i\to\infty}\frac{\lambda_{1}}{R}(p_{i})=0,$
contradicting (3.5). Thus, Case 2 is also impossible.
Combining both cases, we conclude that $M^{-}=\emptyset$. Therefore, $(M^{n},g,f)$ has nonnegative Ricci curvature. ∎
Part II: $(M^{n},g,f)$ has positive curvature operator $Rm>0$.
First of all, when the Weyl tensor $W$ vanishes, we may express the smallest sectional curvature as (see, e.g., [89, Proposition 3.1])
$\mu:=\min_{i,j}R_{ijij}=R_{1212}=\frac{1}{n-2}\big[\lambda_{1}+\lambda_{2}-\frac{R}{n-1}\big],$
where, as before, $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$ are the Ricci eigenvalues.
Remark 3.1.
We also note that, for any manifold with vanishing Weyl tensor, $\mu$ also coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [89, Proposition 3.1(i)]). In particular, when $W=0$, nonnegative sectional curvature is equivalent to nonnegative curvature operator $Rm\geq 0$.
Next, we derive the following elliptic differential inequality for the smallest sectional curvature in the barrier sense.
Lemma 3.2.
Let $(M^{n},g,f)$, $n\geq 3$, be an $n$-dimensional gradient Ricci soliton with vanishing Weyl tensor $W\equiv 0$ and positive scalar curvature $R>0$. Then, in the barrier sense, we have
$(n-2)\Delta_{F}\left(\frac{\mu}{R}\right)=\Delta_{F}\left(\frac{\lambda_{1}+\lambda_{2}}{R}\right)\leq\frac{2E}{(n-1)(n-2)R^{2}},\qquad(F=f-2\log R)$
where
$\begin{split}E&=-(n-2)\lambda_{1}^{2}\sum_{j=3}^{n}(\lambda_{j}-\lambda_{1})-(n-2)\lambda_{2}^{2}\sum_{j=3}^{n}(\lambda_{j}-\lambda_{2})\\ &\quad-\bigl(R-(n-2)\lambda_{1}\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 1\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}-\Biggl(\sum_{j=3}^{n}\bigl(\lambda_{j}-\lambda_{2}\bigr)\Biggr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &\quad-(\lambda_{1}+\lambda_{2})\sum_{j=3}^{n}(\lambda_{2}+\lambda_{j}-2\lambda_{1})(\lambda_{j}-\lambda_{2})-(\lambda_{1}+\lambda_{2})\sum_{\begin{subarray}{c}i>j\geq 3\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}.\end{split}$
Moreover, if the Ricci curvature is $2$-nonnegative then $E\leq 0$.
Proof.
First of all, since $\mu=\frac{1}{n-2}\big(\lambda_{1}+\lambda_{2}-\frac{R}{n-1}\big)$, we have
$(n-2)\Delta_{F}\left(\frac{\mu}{R}\right)=\Delta_{F}\left(\frac{\lambda}{R}\right),$
where $F=f-2\log R$ and $\lambda=\lambda_{1}+\lambda_{2}$.
Next, we proceed to compute $\Delta_{F}(\lambda/R)$. To begin with, for any fixed point $p_{0}\in M$, let $\{e_{i}\}$ be an orthonormal basis that diagonalizes the Ricci tensor so that $Rc(e_{i})=\lambda_{i}e_{i}$ with $\lambda_{1}\leq\cdots\leq\lambda_{n}$. We then extend $\{e_{i}\}$ to a local orthonormal frame in the neighborhood $U$ of $p_{0}$ by parallel transport along radial geodesics emanating from $p_{0}$. The resulting local orthonormal frame in $U$, denoted by $\{\tilde{e}_{i}\}$, satisfy
$\nabla\tilde{e}_{i}(p_{0})=\Delta\tilde{e}_{i}(p_{0})=0.$ (3.6)
Now we define
$\tilde{\lambda}(p)=\big(Rc(\tilde{e}_{1},\tilde{e}_{1})+Rc(\tilde{e}_{2},\tilde{e}_{2})\big)(p).$
On the other hand, for any $p\in M$, it is clear that
$(\lambda_{1}+\lambda_{2})(p)=\inf\Big\{Rc(\tau_{1},\tau_{1})+Rc(\tau_{2},\tau_{2})\Big|\{\tau_{i}\}\text{ is an orthonormal basis of }T_{p}M\Big\}.$
Consequently, at any $p\in U$, we have
$\tilde{\lambda}(p)\geq\lambda(p)=\lambda_{1}(p)+\lambda_{2}(p),$
with equality at $p_{0}$, i.e., $\tilde{\lambda}$ is a barrier function for $\lambda$.
Evaluating at $p_{0}$ in the barrier sense, using (3.6) and [54, (2.8)] (see also [80, Lemma 2.5]), we obtain
$\begin{split}\Delta_{f}\lambda&=\Delta_{f}(\lambda_{1}+\lambda_{2})\\ &\leq\Delta_{f}\big(Rc(\tilde{e}_{1},\tilde{e}_{1})+Rc(\tilde{e}_{2},\tilde{e}_{2})\big)\\ &=(\Delta_{f}Rc)(\tilde{e}_{1},\tilde{e}_{1})+(\Delta_{f}Rc)(\tilde{e}_{2},\tilde{e}_{2})\\ &=2\rho\lambda_{1}-\frac{2nR}{(n-1)(n-2)}\lambda_{1}+\frac{4}{n-2}\lambda_{1}^{2}-\frac{2}{n-2}\left(|Rc|^{2}-\frac{R^{2}}{n-1}\right)\\ &\quad+2\rho\lambda_{2}-\frac{2nR}{(n-1)(n-2)}\lambda_{2}+\frac{4}{n-2}\lambda_{2}^{2}-\frac{2}{n-2}\left(|Rc|^{2}-\frac{R^{2}}{n-1}\right).\end{split}$
Let $F=f-2\log R$. By direct computations as in the proof of Lemma 3.1 and using the above equation, we obtain
$\begin{split}\Delta_{F}\left(\frac{\lambda}{R}\right)&=\frac{1}{R}\Delta_{f}\lambda-\frac{\lambda}{R^{2}}\Delta_{f}R\\ &=\frac{1}{R}\Delta_{f}(\lambda_{1}+\lambda_{2})-\frac{\lambda_{1}+\lambda_{2}}{R^{2}}\Delta_{f}R\\ &\leq\frac{2E}{(n-1)(n-2)R^{2}},\end{split}$
where $E=h_{1}+h_{2}$ and, for $i=1,2$,
$h_{i}=(n-2)\lambda_{i}^{2}(n\lambda_{i}-R)+\bigl((n-2)\lambda_{i}-R\bigr)\Biggl((n-1)\sum_{j\neq i}^{n}\lambda_{j}^{2}-\Bigl(\sum_{j\neq i}^{n}\lambda_{j}\Bigr)^{2}\Biggr).$
It remains to express $E$ in the form as given in Lemma 3.2. For this purpose, we rewrite
$n\lambda_{1}-R=-(\lambda_{2}-\lambda_{1})-\sum_{j=3}^{n}(\lambda_{j}-\lambda_{1}),$
$n\lambda_{2}-R=(\lambda_{2}-\lambda_{1})-\sum_{j=3}^{n}(\lambda_{j}-\lambda_{2}),$
and observe that
$(n-1)\sum_{j\neq 1}^{n}\lambda_{j}^{2}-\Bigl(\sum_{j\neq 1}^{n}\lambda_{j}\Bigr)^{2}=\sum_{\begin{subarray}{c}i>j\\ i,j\neq 1\end{subarray}}(\lambda_{i}-\lambda_{j})^{2},$
$(n-1)\sum_{j\neq 2}^{n}\lambda_{j}^{2}-\Bigl(\sum_{j\neq 2}^{n}\lambda_{j}\Bigr)^{2}=\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}.$
By substituting the above identities into $E=h_{1}+h_{2}$, we get
$\begin{split}E&=(n-2)(\lambda_{2}-\lambda_{1})^{2}(\lambda_{1}+\lambda_{2})-(n-2)\lambda_{1}^{2}\sum_{j=3}^{n}(\lambda_{j}-\lambda_{1})\\ &\quad-(n-2)\lambda_{2}^{2}\sum_{j=3}^{n}(\lambda_{j}-\lambda_{2})+\bigl((n-2)\lambda_{1}-R\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 1\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &\quad+\bigl((n-2)\lambda_{2}-R\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}.\end{split}$ (3.7)
One can notice that every term, except possibly the first one and the last one, in the above expression (3.7) of $E$ is nonpositive. On the other hand, the last term can be rewritten as
$\begin{split}\bigl((n-2)\lambda_{2}-R\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}&=-\Bigg(\lambda_{1}+\lambda_{2}+\sum_{j=3}^{n}(\lambda_{j}-\lambda_{2})\Bigg)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &=-(\lambda_{1}+\lambda_{2})\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &\quad-\Biggl(\sum_{k=3}^{n}(\lambda_{k}-\lambda_{2})\Biggr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\end{split}$ (3.8)
Combining (3.7) and (3.8), we arrive at the desired expression of $E$. Also, it is clear $E\leq 0$ when $\lambda_{1}+\lambda_{2}\geq 0$, i.e., the Ricci curvature is 2-nonnegative. This finishes the proof of the lemma. ∎
Now, we are ready to complete the proof of Theorem 3.1, i.e., $(M^{n},g,f)$ has positive curvature operator $Rm>0$.
Proof of $Rm>0$.
First, we claim that $(M^{n},g,f)$ must have nonnegative sectional curvature. Suppose not. Then, as before, we argue by contradiction. Assume the claim fails. Then the set
$M^{-}:=\{p\in M:\mu(p)<0\}$
is nonempty, where
$\mu=\min_{i,j}R_{ijij}=R_{1212}=\frac{1}{n-2}\big[\lambda_{1}+\lambda_{2}-\frac{R}{n-1}\big]$
is the smallest sectional curvature. Moreover,
$\epsilon_{0}:=\inf_{M}\left(\frac{\mu}{R}\right)<0.$ (3.9)
Case 1: Interior infimum. Suppose the negative infimum $\epsilon_{0}$ is attained at some point $p_{0}\in M$. By Theorem 3.1 (i), we know that the Ricci curvature is nonnegative. Then, by Lemma 3.2, we have
$(n-2)\Delta_{F}\left(\frac{\mu}{R}\right)=\frac{2E}{(n-1)(n-2)R^{2}}\leq 0.\qquad(F=f-2\log R)$
By Calabi’s barrier maximum principle (Lemma 2.2), the ratio $\mu/R$ must be constant on some bounded connected domain $p_{0}\in\Omega\subset M^{-}$. Thus, we have $E\equiv 0$ on $\Omega$ which, by the expression of $E$ in Lemma 3.2, forces
$\begin{split}(n-2)\lambda_{2}^{2}\sum_{j=3}^{n}(\lambda_{2}-\lambda_{j})&\equiv 0,\\[3.00003pt] (n-2)\lambda_{1}^{2}\sum_{j=3}^{n}(\lambda_{1}-\lambda_{j})&\equiv 0,\\[3.00003pt] \bigl((n-2)\lambda_{1}-R\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 1\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}&\equiv 0.\end{split}$ (3.10)
From the first identity in (3.10), either $\lambda_{2}=0$ or $\lambda_{2}=\lambda_{j}$ for all $j\geq 3$.
$\bullet$ Subcase 1a: $\lambda_{2}=0$ at $p_{0}$. By Theorem 3.1 (i), we have $\operatorname{Ric}\geq 0$, which implies $\lambda_{1}(p_{0})=0$. Then, the third relation in (3.10) yields $\lambda_{i}=\lambda_{j}$ at $p_{0}$ for all $i,j\geq 2$, hence $\lambda_{i}(p_{0})=0$ for all $i$, contradicting the assumption that the scalar curvature $R$ is positive.
$\bullet$ Subcase 1b: $\lambda_{2}=\lambda_{j}$, at $p_{0}$, for all $j\geq 3$. The second relation in (3.10) implies, at $p_{0}$, either $\lambda_{1}=0$ or $\lambda_{1}=\lambda_{j}$ for $j\geq 3$. If $\lambda_{1}=\lambda_{j}$ at $p_{0}$ for all $j\geq 3$, then
$\lambda_{1}(p_{0})=\lambda_{2}(p_{0})=\lambda_{j}(p_{0})\qquad(j\geq 3).$
So at $p_{0}$, we have $0>\lambda_{1}(p_{0})=\cdots=\lambda_{n}(p_{0})$ contradicting $R>0$. Suppose instead that $\lambda_{1}(p_{0})=0$ and $\lambda_{2}(p_{0})=\lambda_{j}(p_{0})$ for all $j\geq 3$. Then, at $p_{0}$, we have
$\mu=\frac{1}{n-2}\big[\lambda_{1}+\lambda_{2}-\frac{R}{(n-1)}\big]=0,$
which is a contradiction to (3.9). Therefore, Case 1 is impossible.
Case 2: Infimum at infinity. Suppose $\mu/R$ attains its negative infimum at infinity. Then, by Lemma 6.2 and essentially the same argument as in the proof of $Rc\geq 0$ in Part I, we can rule out Case 2. We omit the details.
Hence, we conclude that $M^{-}=\emptyset$. Therefore, $M$ has nonnegative sectional curvature. As nonnegative sectional curvature is equivalent to nonnegative curvature operator for manifolds with vanishing Weyl tensor $W=0$, it follows that the expanding soliton $(M,g,f)$ has nonnegative curvature operator $Rm\geq 0$. In fact, $\mu$ coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [89, Proposition 3.1(i)]).
Finally, it remains to prove $(M,g,f)$ has positive curvature operator $Rm>0$. To begin with, following the proof of [25, Theorem 1.3], we observe that the null space of the curvature operator, $\ker(Rm)$, is invariant under parallel translation. Indeed, consider the canonical immortal solution
$g(t)=(1+t)\Phi(t)^{\ast}(g),\quad-1<t<\infty,$
to the Ricci flow induced by the expanding soliton $(M^{n},g,f)$, with $g(0)=g$ for $t\in[0,1]$. By the evolution equation of the curvature operator (see [60])
$\partial_{t}Rm(t)=\Delta Rm(t)+2\big(Rm^{2}(t)+Rm^{\sharp}(t)\big)$
and the fact that $(M,g)$ has nonnegative curvature operator $Rm\geq 0$, it follows that the quadratic term $Rm(t)^{2}+Rm(t)^{\sharp}\geq 0$ for all $t\in[0,1]$. Thus, by Hamilton’s strong maximum principle (see, e.g., [27, Theorem 2.2.1]), there exists an interval $0<t<\delta$ over which the rank of $Rm(t)$ is constant, and $\ker(Rm(t))$ is invariant under parallel translation. Since $g(t)=(1+t)\Phi(t)^{\ast}(g)$ with $g(0)=g$, we conclude that the rank of $Rm=Rm(0)$ is locally constant and that $\ker(Rm)=\ker(Rm(0))$ is invariant under parallel translation.
Claim. The expanding Ricci soliton $(M^{n},g,f)$ has positive curvature operator.
Proof of Claim. As in[87], we argue by contradiction. Suppose that $M$ does not have positive curvature operator. Since $M$ has nonnegative curvature operator, there exists a point $p_{0}\in M$ such that the smallest eigenvalue of the curvature operator $\mu(p_{0})=0$. Because the rank of the curvature operator $Rm$ is locally constant, it follows that $\mu\equiv 0$ on a neighborhood $U$ of $p_{0}$. Moreover, as $\ker(Rm)$ is invariant under parallel translation, there exists an orthogonal decomposition of the tangent bundle $TU=V_{1}\oplus V_{2}$, where $V_{1}=\ker(Rm)$ with $\dim(V_{1})\geq 1$, and both $V_{1}$ and $V_{2}$ are invariant under parallel translation. By [60, Lemma 9.1], $M$ locally splits off at least one Euclidean factor. Since expanding Ricci solitons are real analytic, completeness implies that this local splitting extends globally. However, this contradicts the assumption that $M$ is asymptotically conical: any Euclidean factor of $M$ would force the asymptotic cone $\mathcal{C}$ to split off a Euclidean factor, thereby violating the regularity assumption that $\mathcal{C}$ has only an isolated singularity at the tip, unless $\mathcal{C}$ is a Euclidean space with flat metric. In this latter case, a flat Euclidean space is not a cone with positive scalar curvature, which is a contradiction. Consequently, $M$ must have positive curvature operator.
This completes the proof of the Claim, and the proof of Theorem 3.1. ∎
Proof of Corollary 1.2.
Let $(M^{4},g,f)$ be a non-compact, half conformally flat (i.e., either $W^{+}=0$ or $W^{-}=0$), asymptotically conical Ricci expander with positive scalar curvature. By essentially the same arguments as in Chen-Wang [39] for gradient shrinking and steady Ricci solitons, one can show that $(M^{4},g,f)$ has vanishing $D$-tensor as introduced in [18, 19]. Thus, by [19, Proposition 3.2], $(M^{4},g,f)$ is locally conformally flat. Therefore, it follows from Theorem 3.1 that $(M^{4},g,f)$ has positive curvature operator $Rm>0$. ∎
Finally, we conclude this section by proving Theorem 1.1.
Proof of Theorem 1.1.
First, note that we can diagonalize the curvature operator $Rm$ with eigenvalues $m_{1}\leq m_{2}\leq m_{3}$. Then, the Ricci tensor $Rc$ is diagonalized with eigenvalues
$(m_{1}+m_{2})\leq(m_{1}+m_{3})\leq(m_{2}+m_{3}),$
and the scalar curvature is given by $R=2(m_{1}+m_{2}+m_{3})$. Moreover, for $n=3$, the differential equation
$\Delta_{f}Rm=2\rho Rm-2(Rm^{2}+Rm^{\sharp}),$
implies ([60])
$\Delta_{f}m_{1}\leq 2\rho m_{1}-2(m_{1}^{2}+m_{2}m_{3}).$
Then, by direct computations, we have
$\begin{split}\Delta_{F}\left(\frac{m_{1}}{R}\right)&=\frac{1}{R}\Delta_{f}m_{1}-\frac{m_{1}}{R^{2}}\Delta_{f}R\quad\quad(F=f-2\log R)\\ &\leq\frac{2}{R^{2}}\left[m_{1}|Rc|^{2}-R(m_{1}^{2}+m_{2}m_{3})\right]\\ &=\frac{4}{R^{2}}\left[(m_{1}-m_{3})m_{2}^{2}+(m_{1}-m_{2})m_{3}^{2}\right]\\ &\leq 0.\end{split}$ (3.11)
Next, we show $m_{1}\geq 0$ (hence $(M^{3},g,f)$ has nonnegative curvature operator) by contradiction. Suppose not. Then, the set
$M^{-}:=\{p\in M:m_{1}(p)<0\}$
is nonempty, and
$\epsilon:=\inf_{M}\left(\frac{m_{1}}{R}\right)<0.$
Case 1: Interior infimum. Suppose the negative infimum $\epsilon$ is attained at some $p_{0}\in M$ such that $m_{1}(p_{0})<0$. By (3.11), we have
$\begin{split}\Delta_{F}\left(\frac{m_{1}}{R}\right)\leq\frac{4}{R^{2}}\left[(m_{1}-m_{3})m_{2}^{2}+(m_{1}-m_{2})m_{3}^{2}\right]\leq 0.\end{split}$
Then, Calabi’s barrier strong maximum principle (Lemma 2.2) implies the ratio $m_{1}/R$ must be constant on a neighborhood of $p_{0}$. In particular, in that neighborhood of $p_{0}$, we have
$(m_{1}-m_{3})m_{2}^{2}\equiv 0,\quad(m_{1}-m_{2})m_{3}^{2}\equiv 0.$ (3.12)
Since $m_{1}(p_{0})<0$ and the scalar curvature $R>0$, the first equation in (3.12) forces $m_{2}(p_{0})=0$. However, the second equation in (3.12) then implies $m_{3}(p_{0})=0$, which contradicts the assumption of positive scalar curvature. Thus, Case 1 is ruled out.
Case 2: Infimum at infinity. Assume instead that $m_{1}/R$ attains its negative infimum at infinity. By Lemma 6.1, any 3-dimensional cone with positive scalar curvature automatically has positive curvature operator/sectional curvature, and the smallest eigenvalue of its curvature operator is zero. Then, as in Part I of the proof of Theorem 3.1, we get a contradiction. Thus, Case 2 is impossible.
Combining both cases, we conclude that $M^{-}=\emptyset$, so $M$ has nonnegative curvature operator or, equivalently, nonnegative sectional curvature (for $n=3$).
Finally, by the same argument as in the proof of $Rm>0$ in Theorem 3.1, we conclude that $M^{3}$ has positive curvature operator. This completes the proof of Theorem 1.1. ∎
Remark 3.2.
Theorem 1.1 also follows from Theorem 1.2, which is valid for $n=3$ since the Weyl tensor vanishes automatically, and Lemma 6.1 on the curvature operator of 3D cones.
4. 4D asymptotically conical Ricci expanders with (half) PIC
In this section, we study curvature pinching in four-dimensional asymptotically conical gradient Ricci expanders with (half) PIC. In particular, we prove Theorem 1.3 and Theorem 1.4.
We first investigate the positivity of the self-dual curvature operator $A$ (respectively, the anti-self-dual curvature operator $C$), as defined in the curvature decomposition (2.2), for asymptotically conical gradient expanding Ricci solitons with $A_{2}\geq 0$ (respectively, $C_{2}\geq 0$). We assume in addition that the asymptotic cone has nonnegative self-dual (or anti-self-dual) curvature operator and positive scalar curvature. Under these conditions, we prove the following stronger result, from which Theorem 1.3 follows.
Theorem 4.1.
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton. Suppose that the asymptotic cone has positive scalar curvature and satisfies either $A\geq 0$ or $C\geq 0$. Then:
  • (a)
    If $(M^{4},g,f)$ satisfies $A_{2}\geq 0$ or $C_{2}\geq 0$, then $A\geq 0$ or $C\geq 0$ on $M^{4}$.
  • (b)
    If $(M^{4},g,f)$ satisfies $A_{2}>0$ or $C_{2}>0$, then $A>0$ or $C>0$ on $M^{4}$.
Proof.
Again, as first mentioned in Section 3, the assumption of the asymptotic cone having positive scalar curvature implies that $(M^{n},g,f)$ itself has positive scalar curvature $R>0$.
(a) Without loss of generality, we may assume $A_{2}\geq 0$. Again, as in the proof of Theorem 1.2, we argue by contradiction. Suppose instead that the claim fails. Then the set
$M^{-}:=\{p\in M:A_{1}(p)<0\}$
is nonempty, and we have
$\epsilon^{\prime}:=\inf_{M}\left(\frac{A_{1}}{R}\right)<0.$
Case 1: Interior infimum. Suppose the negative infimum $\epsilon^{\prime}$ is attained at some point $p_{0}\in M$. Then, there exists a neighborhood $\Omega\ni p_{0}$, such that $A_{1}(p)<0$ for all $p\in\Omega$. Thus, by Lemma 2.1 and direct computation, for $F=f-2\log R$, we have
$\Delta_{F}\left(\frac{A_{1}}{R}\right)\leq\frac{2}{R^{2}}\left[A_{1}|Rc|^{2}-R(A_{1}^{2}+B_{1}^{2}+2A_{3}A_{2})\right]\leq 0,$
in the barrier sense on $\Omega$.
Now, by Calabi’s barrier strong maximum principle (Lemma 2.2), the ratio $A_{1}/R$ must be constant on $\Omega$. In particular, we have
$A_{1}|Rc|^{2}-R(A_{1}^{2}+B_{1}^{2}+2A_{3}A_{2})\equiv 0,$
which implies $A_{1}|Rc|^{2}=R(A_{1}^{2}+B_{1}^{2}+2A_{3}A_{2})\equiv 0$ on $\Omega$.
Since $A_{1}/R$ attains its negative infimum at $p_{0}$, $A_{1}(p_{0})\neq 0$ so we must have $|Rc|^{2}=0$. then, the scalar curvature is zero. This contradicts the fact that the scalar curvature is positive. Hence, Case 1 is ruled out.
Case 2: Infimum at infinity. If $A_{1}/R$ attains its negative infimum at infinity, then by the assumptions on the asymptotic cone and the same argument in the proof of Theorem 3.1, we can rule out Case 2.
Combining both cases, we conclude that $M^{-}=\emptyset$. Therefore, $M$ satisfies $A\geq 0$, completing the proof of part (a).
(b) Without loss of generality, we assume $A_{2}>0$. By part (a), this implies $A_{1}\geq 0$. Following [24, Proposition 3.1(b)], we prove that $A_{1}>0$ by contradiction. Suppose $A_{1}(p_{0})=0$ at some point $p_{0}\in M^{4}$. Then $A_{1}$ attains its minimum at $p_{0}$. Let $\eta\in\wedge_{p_{0}}^{+}(M)$ be a null eigenvector of $A$ such that $A(\eta,\eta)=A_{1}(p_{0})=0$ at $p_{0}$. Extend $\eta$ to a local section (still denoted by $\eta$) by parallel transport along geodesics emanating from $p_{0}$.
At $p_{0}$, in the barrier sense, Lemma 2.1 yields
$\begin{split}0&\leq\Delta_{f}A_{1}\\ &\leq\Delta_{f}A(\eta,\eta)\\ &=(\Delta_{f}A)(\eta,\eta)\\ &\leq 2\bigl(\rho A_{1}-A_{1}^{2}-2A_{2}A_{3}-B_{1}^{2}\bigr)\\ &<0,\end{split}$
where we used the assumption $A_{3}\geq A_{2}>0$ in the last inequality. This contradiction shows that $A_{1}>0$ on $M^{4}$.
This completes the proof of Theorem 4.1. ∎
Next, we prove Theorem 1.4.
Proof of Theorem 1.4.
Let $0\leq B_{1}\leq B_{2}\leq B_{3}$ be the singular eigenvalues of the matrix $B$ and ${\lambda}_{1}\leq{\lambda}_{2}\leq{\lambda}_{3}\leq{\lambda}_{4}$ be the eigenvalues of the Ricci tensor Rc{Rc}. Then, by [25, Lemma 2.2], the sum of the least two eigenvalues of $Rc$ is given by
$\lambda_{1}+\lambda_{2}=\tfrac{1}{2}(R-4B_{3}).$
Thus, 2-nonnegative Ricci curvature is equivalent to $u:=R-4B_{3}\geq 0.$
Now, by Lemma 2.1 and essentially the same computations as in the proof of [25, Theorem 3.1(a)], we have
$\begin{split}\Delta_{f}u&=\Delta(R-4B_{3})\\ &\leq 2\rho u-\left[2|Rc|^{2}-8(A_{3}B_{3}+C_{3}B_{3}+2B_{1}B_{2})\right]\\ &\leq 2\rho u-8(A_{2}+A_{1}+C_{2}+C_{1})B_{3}.\end{split}$
Then, by Lemma 2.1 and direct computations, for $F=f-2\log R$, we have
$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&=\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{2}{R^{2}}\big[u|Rc|^{2}-4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\big],\end{split}$ (4.1)
Next, we prove, again by contradiction, that $(M^{4},g,f)$ has 2-nonnegative Ricci curvature. Suppose instead that the Ricci curvature is not 2-nonnegative. Then,
$M^{-}:=\{p\in M:u(p)<0\}$
is nonempty, and we have
$\tilde{\epsilon}:=\inf_{M}\left(\frac{u}{R}\right)<0.$ (4.2)
Case 1: Interior infimum. Suppose the negative infimum $\tilde{\epsilon}$ is attained at some point $p_{0}\in M$. Then, there exists a neighborhood $\Omega\ni p_{0}$, such that $u<0$ on $\Omega$. By (4.1) and the assumption that $(M^{4},g,f)$ has PIC, so that $A_{1}+A_{2}>0$ and $C_{1}+C_{2}>0$, we have
$\Delta_{F}\left(\frac{u}{R}\right)\leq\frac{2}{R^{2}}\left[u|Rc|^{2}-4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\right]\leq 0$
on $\Omega$, where we have used the fact that $B_{3}\geq 0$. Thus, by Calabi’s barrier strong maximum principle (Lemma 2.2), the ratio $u/R$ must be constant on $\Omega$. In particular, we have
$u|Rc|^{2}-4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\equiv 0,$
which implies $u|Rc|^{2}=4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}=0$ on $\Omega$.
Since $u(p_{0})<0$ by (4.2), we must have $|Rc|^{2}=0$ at $p_{0}$. But this is a contradiction to the scalar curvature $R>0$. Hence, Case 1 is ruled out.
Case 2: Infimum at infinity. Suppose that $u/R$ attains its negative infimum at infinity. Since $(M^{4},g,f)$ has PIC, the asymptotic cone must have WPIC. By Lemma 6.3, the Ricci curvature of the asymptotic cone is therefore nonnegative. Then by similar arguments in the proof of Theorem 3.1, we can rule out Case 2.
Combining both cases, we conclude that $M^{-}=\emptyset$. Therefore, $M$ satisfies $u\geq 0$, completing the proof of 2-nonnegativity of the Ricci curvature.
Finally, given that $(M^{4},g,f)$ has 2-nonnegative Ricci curvature, following the proof in [25, Theorem 3.1(b)], we shall prove 2-positive Ricci curvature by contradiction. We consider the quadratic form $Z:=RI-4\sqrt{B{}^{t}B}$, where $I$ is the 3 by 3 identity matrix. By the 2-nonnegativity of the Ricci curvature, we know that $Z\geq 0$ and that 2-positive Ricci curvature is equivalent to $Z>0$. Now, we denote the eigenvalues of $Z$ by
$0\leq Z_{1}\leq Z_{2}\leq Z_{3}.$
Assume that $Z$ has a null eigenvector at some point $p_{0}$. Then $Z_{1}$ attains its minimum at $p_{0}$. Let $\eta\in\wedge_{p_{0}}^{+}(M)$ be a null eigenvector of $Z$ such that $Z(\eta,\eta)=Z_{1}(p_{0})=0$ at $p_{0}$. Extend $\eta$ to a local section (still denoted by $\eta$) by parallel transport along geodesics emanating from $p_{0}$. Then, at $p_{0}$, in the barrier sense, we have
$\begin{split}0&\leq\Delta_{f}Z_{1}\\ &\leq\Delta_{f}Z(\eta,\eta)\\ &=(\Delta_{f}Z)(\eta,\eta)\\ &\leq 2\rho Z_{1}-8(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\\ &<0,\end{split}$
where we have used the PIC condition and $B_{3}(p_{0})=\tfrac{R}{4}(p_{0})>0$ in the last inequality. Thus, we get a contradiction. Therefore, the Ricci curvature is 2-positive.
This completes the proof of Theorem 1.4. ∎
Finally, by Lemma 6.3 and an argument similar to that used in the proof of Theorem 1.4, we obtain the following slightly stronger result than Theorem 1.4.
Theorem 4.2.
Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton. Suppose that the asymptotic cone has positive scalar curvature, and satisfies $A\geq 0$ and $C\geq 0$. Then:
  • (a)
    If $(M^{4},g,f)$ satisfies $A_{2}\geq 0$ and $C_{2}\geq 0$, then the Ricci curvature is $2$-nonnegative.
  • (b)
    If $(M^{4},g,f)$ satisfies $A_{2}>0$ and $C_{2}>0$, then the Ricci curvature is $2$-positive.
5. General lemma and further applications
In this section, we formulate a fairly general method that can be effectively applied to prove generalized Hamilton-Ivey type curvature pinching estimates for a class of non-compact, asymptotically conical gradient Ricci solitons. As applications, we obtain several analogues of other known curvature pinching results for ancient solutions, in the setting of asymptotically conical expanding Ricci solitons–including Theorem 1.5 as stated in the introduction, and Theorem 5.1 below.
5.1. A general lemma
By identifying common patterns in how we have proved the curvature pinching theorems in Section 4, we are led to the following
Lemma 5.1.
Let $(M^{n},g,f)$ be an $n$-dimensional non-compact gradient expanding Ricci soliton, satisfying Eq. (1.1) and with positive scalar curvature. Suppose $u:M^{n}\rightarrow\mathbb{R}$ is a Lipschitz function and satisfies the differential inequality
$\Delta_{f}u\leq 2\rho u-v$ (5.1)
in the barrier sense, where $v\geq 0$ is a nonnegative function on $M^{n}$.
  • (i)
    If for any sequence of points $\{p_{i}\}\subset M^{n}$, with $p_{i}\rightarrow\infty$, we have
    $\liminf_{i\rightarrow\infty}\left(\frac{u}{R}\right)\geq 0.$ (5.2)
    Then, $u\geq 0$ on $M^{n}$.
  • (ii)
    If in addition $v>0$, then $u>0$ on $M^{n}$.
Proof.
(i) First of all, we shall compute the differential equation of $u/R$ in the barrier sense. By direct computations, we have
$\begin{split}\Delta_{f}\left(\frac{u}{R}\right)=\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R-\frac{2}{R^{2}}\langle\nabla u,\nabla R\rangle+\frac{2u}{R^{3}}|\nabla R|^{2}.\end{split}$ (5.3)
Let $F=f-2\log R$. Substituting (5.1) and the formula for $\Delta_{f}R$ in Lemma 2.1 into (5.3), we obtain
$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&\leq\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{1}{R}(2\rho u-v)-\frac{u}{R^{2}}(2\rho R-2|Rc|^{2})\\ &\leq\frac{1}{R^{2}}(2u|Rc|^{2}-Rv).\end{split}$ (5.4)
Now, we prove by contradiction again. Suppose the lemma fails. Then the set
$M^{-}:=\{p\in M:u(p)<0\}$
is nonempty, and
$\delta:=\inf_{M}\left(\frac{u}{R}\right)<0.$
Case 1: Interior infimum. Suppose the negative infimum $\delta$ is attained at some $p_{0}\in M$. Then, in a neighborhood $\Omega\ni p_{0}$, we have $u<0$ on $\Omega$. Thus, by (5.4), we have
$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&\leq\frac{1}{R^{2}}(2u|Rc|^{2}-Rv)\leq 0\end{split}$
in the barrier sense on $\Omega$. By Calabi’s barrier strong maximum principle (Lemma 2.2), the ratio $u/R$ must be constant on $\Omega$. Since $R>0$, $v\geq 0$ and $u<0$ on $\Omega$, it follows that
$u|Rc|^{2}\equiv 0,$
which forces $|Rc|=0$ on $\Omega$, a contradiction to $R>0$. Thus, Case 1 is ruled out.
Case 2: Infimum at infinity. Assume instead that $u/R$ attains its negative infimum at infinity. Then there exists a sequence $\{p_{i}\}\subset M$ with $p_{i}\to\infty$ such that
$\lim_{i\to\infty}\frac{u}{R}(p_{i})=\delta<0.$
However, this contradicts our assumption (5.2). Thus, Case 2 is impossible.
Combining both cases, we conclude that $M^{-}=\emptyset$, so $u\geq 0$ on $M^{n}$.
(ii) We prove $u>0$ by contradiction. Suppose $u(p_{0})=0$ at some point $p_{0}\in M$. Then, since $u\geq 0$ on $M^{n}$, it follows that $u$ attains its minimum at $p_{0}$. Then, at $p_{0}$, in the barrier sense
$\begin{split}0\leq\Delta_{f}u\leq 2\rho u-v<0,\end{split}$
where we have used the assumption $v>0$. This is a contradiction. ∎
Remark 5.1.
In particular, if an expanding Ricci soliton $(M^{n},g,f)$ is asymptotically conical, and if its asymptotic cone has positive scalar curvature and satisfies $u\geq 0$, then $(M^{n},g,f)$ fulfills the asymptotic condition (5.2) in Lemma 5.1.
Remark 5.2.
Lemma 5.1 also applies to gradient shrinking and steady Ricci solitons.
5.2. The proof of Theorem 1.5
In this subsection, we apply Lemma 5.1 to prove Theorem 1.5. Recall that, by uniformly PIC we mean that $(M^{4},g)$ has PIC and satisfies in addition the pointwise pinching condition
$\max\{A_{3},B_{3},C_{3}\}\leq\Lambda\min\{A_{1}+A_{2},C_{1}+C_{2}\},$
for some constant $\Lambda\geq 1$.
Proof of Theorem 1.5.
First of all, it is easy to see that the inequality $B_{3}^{2}\leq A_{1}C_{1}$ implies nonnegative curvature operator $Rm\geq 0$ for $(M^{4},g)$; see, e.g., [43, Lemma 4.4] for a proof. Hence, to prove Theorem 1.5, it suffices to establish the inequality $B_{3}^{2}\leq A_{1}C_{1}$.
We shall prove the inequality $B_{3}^{2}\leq A_{1}C_{1}$ in three steps as in [7]. The main computations in each step below essentially come from Brendle’s work [7], in which he used the pinching estimates of Hamilton [63] to show that a gradient steady Ricci soliton with UPIC must have positive curvature operator.
Step 1. To show $A_{3}\leq(6\Lambda^{2}+1)A_{1}$ and $C_{3}\leq(6\Lambda^{2}+1)C_{1}.$
By the same computations as in [7, Lemma 6.1], we have
$\Delta_{f}[(6\Lambda^{2}+1)A_{1}-A_{3}]\leq 2\rho[(6\Lambda^{2}+1)A_{1}-A_{3}]-A_{3}^{2}.$
Moreover, since the asymptotic cone $\mathcal{C}$ is a non-flat Euclidean cone, by (6.8) and (6.9), on the cone $\mathcal{C}$, we have
$\bar{A}_{i}=\bar{C}_{j}=\bar{B}_{k},\quad 1\leq i,j,k\leq 3,$
where the bar denotes the corresponding curvature quantities on the asymptotic cone $\mathcal{C}$. Hence, the inequality $A_{3}\leq(6\Lambda^{2}+1)A_{1}$ follows from Lemma 5.1 with $u=(6\Lambda^{2}+1)A_{1}-A_{3}$. Similarly, we have $C_{3}\leq(6\Lambda^{2}+1)C_{1}.$
Step 2. To show $4B_{3}^{2}\leq(A_{1}+A_{2})(C_{1}+C_{2}).$
Following [7], we prove by contradiction. Suppose that
$\gamma=\sup_{M}\frac{2B_{3}}{\sqrt{(A_{1}+A_{2})(C_{1}+C_{2})}}>1.$
Let $w_{1}:=\tfrac{1}{2}\sqrt{(A_{1}+A_{2})(C_{1}+C_{2})}$. By the same computations as in [7, Lemma 6.2], we can find a positive constant $\delta_{1}>0$ such that
$\Delta_{f}(\gamma w_{1}-B_{3}-\delta_{1}R)\leq 2\rho(\gamma w_{1}-B_{3}-\delta_{1}R)-\delta_{1}|Rc|^{2}.$
On the other hand, since $\gamma>1$, on the asymptotic non-flat Euclidean cone $\mathcal{C}$ we can choose $\bar{\delta}_{1}>0$ small enough such that
$\gamma\bar{w}_{1}-\bar{B}_{3}-\bar{\delta}_{1}\bar{R}>0.$
Hence, by Lemma 5.1 with $u=\gamma w_{1}-B_{3}-\delta_{1}R$, we obtain $\gamma w_{1}-B_{3}-\delta_{1}R\geq 0$, which contradicts the definition of $\gamma$. Therefore, $\gamma\leq 1$, completing the proof of Step 2.
Step 3. To show $B_{3}^{2}\leq A_{1}C_{1}.$
Again, following [7], we argue by contradiction. Suppose that
$\gamma^{\prime}=\sup_{M}\frac{B_{3}}{\sqrt{A_{1}C_{1}}}>1.$
Let $w_{2}:=\sqrt{A_{1}C_{1}}$. Then by the same computations as in [7, Proposition 6.3], we can find a positive constant $\delta_{2}>0$ such that
$\Delta_{f}(\gamma^{\prime}w_{2}-B_{3}-\delta_{2}R)\leq 2\rho(\gamma^{\prime}w_{2}-B_{3}-\delta_{2}R)-\delta_{2}|Rc|^{2}.$
On the other hand, since $\gamma^{\prime}>1$, on the asymptotic non-flat Euclidean cone $\mathcal{C}$ we can choose $\bar{\delta}_{2}>0$ small enough such that
$\gamma^{\prime}\bar{w}_{2}-\bar{B}_{3}-\bar{\delta}_{2}\bar{R}>0.$
Hence, as in the proof of Step 2, Step 3 follows from Lemma 5.1 with $u=\gamma^{\prime}{w}_{2}-{B}_{3}-{\delta}_{2}{R}$. Therefore, $(M^{4},g,f)$ has nonnegative curvature operator.
Finally, by applying the same argument as in the proof of Theorem 3.1 for showing $Rm>0$, it follows that $(M^{4},g,f)$ has positive curvature operator. This concludes the proof of Theorem 1.5. ∎
Remark 5.3.
The assumption of the asymptotic cone being a (non-flat) Euclidean cone is used in the proofs of Step 2 and Step 3. Indeed, by (6.8) and (6.9), requiring either $4B_{3}^{2}\leq(A_{1}+A_{2})(C_{1}+C_{2})$ or $B_{3}^{2}\leq A_{1}C_{1}$ to hold on the asymptotic cone forces the link of the cone to be a (spherical) space form.
5.3. Additional curvature pinching results
Furthermore, applying Lemma 5.1, we obtain several additional curvature pinching results for asymptotically conical expanding Ricci solitons, which are analogous to results previously established for ancient solutions by Li-Wang [72], Bamler-Cabezas-Rivas-Wilking [3], Li-Ni [70], Li [71], Cho-Li [43], Chen [40], and others.
Theorem 5.1.
Let $(M^{n},g,f)$ be an $n$-dimensional non-compact, asymptotically conical, gradient expanding Ricci soliton.
  • (a)
    Suppose $(M^{n},g,f)$ has $2$-nonnegative curvature operator and the asymptotic cone has positive scalar curvature. Then, $(M^{n},g,f)$ must have positive curvature operator.
  • (b)
    Suppose $(M^{n},g,f)$ has WPIC1 and that the asymptotic cone has positive scalar curvature and WPIC2. Then, $(M^{n},g,f)$ must have WPIC2.
  • (c)
    Suppose $(M^{n},g,f)$, $n\geq 5$, has WPIC and that the asymptotic cone has positive scalar curvature and $2$-nonnegative Ricci curvature. Then, $(M^{n},g,f)$ must have $2$-nonnegative Ricci curvature.
  • (d)
    Suppose $(M^{n},g,f)$ is Kähler and has nonnegative orthogonal bisectional curvature. If the asymptotic cone has positive scalar curvature and WPIC2, then $(M^{n},g,f)$ must have WPIC2.
  • (e)
    Suppose $(M^{n},g,f)$, $n\geq 9$, has UPIC. If the asymptotic cone has positive scalar curvature and WPIC2, then $(M^{n},g,f)$ must have WPIC2.
Sketch of Proof.
It suffices to verify, in each case, that the corresponding least curvature eigenvalue satisfies the differential equation (5.1) in Lemma 5.1.
(a) By the same computations as in [72, Theorem 27] (see also [38, Lemma 2.4]), the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the Riemann curvature operator is satisfied.
(b) By the same computations as in [3, Lemma 4.2], the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the complex sectional curvature is satisfied.
(c) By the same computations as in [70, Proposition 4.2], the differential inequality (5.1) in Lemma 5.1 for the sum of the two least eigenvalues of the Ricci tensor is satisfied.
(d) First of all, by the same computations as in [70, Lemma 6.1], the differential inequality (5.1) in Lemma 5.1 for the least bisectional curvature is satisfied. Hence, by Lemma 5.1, the expanding Kähler-Ricci soliton has nonnegative bisectional curvature. Moreover, by the same arguments as in [71, Theorem 3.3], one can show that the smallest eigenvalue of the complex sectional curvature also satisfies the differential inequality (5.1) in Lemma 5.1.
(e) Finally, by the same computations as in [43, Theorem 3.2] for $n\geq 12$ and in [40] for $9\leq n\leq 11$, the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the complex sectional curvature is satisfied. ∎
Remark 5.4.
Other results that have been derived using B.-L. Chen’s lemma (see, e.g., [43, Corollary 2.4] or [25, Lemma 2.6]) for ancient solutions can similarly be extended to the setting of asymptotically conical gradient expanding Ricci solitons, following the same approach as above.
6. Appendix
In this appendix, we examine the elementary relations between the curvature tensor of an $n$-dimensional cone, $n\geq 3$, with vanishing Weyl tensor and the curvature tensor of its link. Moreover, we analyze the curvature operator decomposition of a $4$-dimensional cone. The resulting facts were used in previous sections.
First of all, let us recall basic curvature relations between a cone and its link. For $n\geq 3$, consider any $n$-dimensional cone
$\mathcal{C}^{n}:=[0,\infty)\times\Sigma^{n-1}$
equipped with the Riemannian metric
$g_{c}=dr^{2}+r^{2}\bar{g},$
where $(\Sigma^{n-1},\bar{g})$ is a closed (n1)(n-1)-dimensional Riemannian manifold. Let $\{\bar{e}_{a}\}_{a\geq 2}$ be a local orthonormal frame of $T\Sigma^{n-1}$. We define
$e_{1}=\partial_{r},\qquad e_{a}=r^{-1}\bar{e}_{a}\quad(a\geq 2),$
so that $\{e_{i}\}_{i\geq 1}$ forms a local orthonormal frame of $T\mathcal{C}$ with respect to $g_{c}$. Then, the Riemann curvature tensor [^3] For the curvature tensor formula of a general warped product space, see O’Neill [78]. $Rm$ of $(\mathcal{C}^{n},g_{c})$ is given by
$R_{1ijk}$ $=0,\quad 1\leq i,j,k\leq n,$ (6.1)
$R_{abcd}$ $=r^{-2}\left[\bar{R}_{abcd}-\big(\bar{g}_{ac}\bar{g}_{bd}-\bar{g}_{ad}\bar{g}_{bc}\big)\right],\quad 2\leq a,b,c,d\leq n,$
where $\bar{R}_{abcd}$ denotes the curvature tensor of the metric $\bar{g}$ of the link $\Sigma^{n-1}$. Moreover, the Ricci tensor $Rc$ of $(\mathcal{C}^{n},g)$ is given by
$R_{1i}$ $=0,\quad 1\leq i\leq n,$ (6.2)
$R_{ab}$ $=r^{-2}\left[\bar{R}_{ab}-(n-2)\bar{g}_{ab}\right],\quad 2\leq a,b\leq n.$
and the scalar curvatures of $(\mathcal{C}^{n},g)$ and $(\Sigma^{n-1},\bar{g})$ are related by
$R=r^{-2}\left[\bar{R}-(n-1)(n-2)\right].$ (6.3)
In addition, the nonzero Weyl curvature tensor $W$ of $(\mathcal{C}^{n},g)$ is given by
$W_{1a1b}$ $=-\frac{1}{(n-2)r^{2}}\left(\bar{R}_{ab}-\frac{\bar{R}}{(n-1)}\bar{g}_{ab}\right),\quad 2\leq a,b\leq n,$ (6.4)
$W_{abcd}$ $=r^{-2}\ \!\overline{W}_{abcd},\quad 2\leq a,b,c,d\leq n.$
6.1. Curvature tensor of cones with vanishing Weyl tensor
First, consider
$\mathcal{C}^{3}=(0,\infty)\times\Sigma^{2},\qquad g_{c}=dr^{2}+r^{2}\bar{g},$
where $(\Sigma^{2},\bar{g})$ is a closed Riemannian surface. Then, the nonzero Riemann curvature tensor components of $(\mathcal{C}^{3},g_{c})$ are given by
$R_{abcd}=r^{-2}\left[\bar{R}_{abcd}-\big(\bar{g}_{ac}\bar{g}_{bd}-\bar{g}_{ad}\bar{g}_{bc}\big)\right],\qquad 2\leq a,b,c,d\leq 3.$
Since $\dim\Sigma=2$, we have
$\bar{R}_{abcd}=\bar{K}\big(\bar{g}_{ac}\bar{g}_{bd}-\bar{g}_{ad}\bar{g}_{bc}\big),$
where $\bar{K}$ is the Gaussian curvature of $(\Sigma^{2},\bar{g})$. Hence
$R_{abcd}=r^{-2}\left(\bar{K}-1\right)\big(\bar{g}_{ac}\bar{g}_{bd}-\bar{g}_{ad}\bar{g}_{bc}\big).$
Therefore, with respect to the basis
$\{e_{1}\wedge e_{2},e_{3}\wedge e_{1},e_{2}\wedge e_{3}\}$
of $\wedge^{2}T\mathcal{C}$, the curvature operator is diagonal with eigenvalues
$R_{1212}=0,\quad R_{1313}=0,\quad R_{2323}=r^{-2}\left(\bar{K}-1\right).$
In summary, we have the following basic fact in dimension $n=3$.
Lemma 6.1.
Let $\mathcal{C}^{3}:=[0,\infty)\times\Sigma^{2}$ be a $3$-dimensional cone equipped with the Riemannian metric $g_{c}=dr^{2}+r^{2}\bar{g}_{\Sigma}$. Then, the only possible nonzero eigenvalue of the curvature operator $Rm$ is the principal sectional curvature
$m:=r^{-2}\left(\bar{K}-1\right),$
where $\bar{K}$ is the Gaussian curvature of $(\Sigma^{2},\bar{g})$. In particular, $(\mathcal{C}^{3},g_{c})$ has nonnegative curvature operator $Rm\geq 0$ if and only if it has nonnegative scalar curvature $R\geq 0$, or equivalently, if and only if $\bar{K}\geq 1$.
Meanwhile, for $n\geq 4$, consider any $n$-dimensional cone
$\mathcal{C}^{n}:=[0,\infty)\times\Sigma^{n-1}$
equipped with the Riemannian metric
$g_{c}=dr^{2}+r^{2}\bar{g},$
where $(\Sigma^{n-1},\bar{g})$ is a closed (n1)(n-1)-dimensional Riemannian manifold. Then, by (6.4), we see that $\mathcal{C}$ has vanishing Weyl curvature $W=0$ if and only if its link $\Sigma$ is a space form. Thus, we immediately have the following
Lemma 6.2.
Let $\mathcal{C}^{n}:=[0,\infty)\times\Sigma^{n-1}$ be an $n$-dimensional ($n\geq 4$) cone with nonnegative scalar curvature, equipped with the Riemannian metric $g_{c}=dr^{2}+r^{2}\bar{g}_{\Sigma}$. Then, $(\operatorname{\mathcal{C}}^{n},g_{c})$ is locally conformally flat but non-flat if and only if the link $(\Sigma^{n-1},\bar{g})$ is a spherical space form, i.e., up to scaling, $(\Sigma^{n-1},\bar{g})$ is isometric to a quotient of the round sphere $\mathbb{S}^{n-1}$. In particular, if $(\operatorname{\mathcal{C}}^{n},g_{c})$ is locally conformally flat and has nonnegative scalar curvature, then it has nonnegative curvature operator $Rm\geq 0$.
6.2. Curvature decomposition and curvature operator of 4D cones
Recall that, with respect to the decomposition
$\wedge^{2}=\wedge^{+}\oplus\wedge^{-}$
on any oriented smooth Riemannian 4-manifold $(M^{4},g)$, the curvature operator of $(M^{4},g)$ admits the following decomposition:
$Rm=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\frac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\frac{R}{12}I\end{pmatrix}$, (6.5)
where $W^{\pm}$ denote the self-dual and anti-self-dual Weyl curvature tensors, respectively, and $\mathring{Rc}$ denotes the traceless Ricci tensor.
We may choose a basis for $\wedge_{p}^{+}$ and for $\wedge_{p}^{-}$ as follows:
[…]
where $\{e_{1},e_{2},e_{3},e_{4}\}$ is any positively oriented orthonormal tangent frame at a point $p$. Here, we have used the metric $g$ to identify the tangent space and the cotangent space at $p$. The inner product on 2-forms is defined by
$\langle X\wedge Y,V\wedge W\rangle=\langle X,V\rangle\langle Y,W\rangle-\langle X,W\rangle\langle Y,V\rangle.$
Observe that, for the matrices $A$ and $C$ in (6.5), we have
[…]
For the matrix $B$, we have
[…]
and
$B_{12}=\tfrac{1}{2}\left(R_{1213}+R_{3413}-R_{1242}-R_{3442}\right)=\tfrac{1}{2}\left(R_{23}-R_{14}\right),\ etc.$
Now, we consider any $4$-dimensional cone
$\mathcal{C}^{4}:=[0,\infty)\times\Sigma^{3}$
equipped with the Riemannian metric
$g_{c}=dr^{2}+r^{2}\bar{g},$
where $(\Sigma^{3},\bar{g})$ is a closed $3$-dimensional Riemannian manifold. On $(\Sigma^{3},\bar{g})$, diagonalize the curvature operator $\overline{Rm}$ with respect to the local 2-frame
$\{\bar{e}_{2}\wedge\bar{e}_{3},\bar{e}_{3}\wedge\bar{e}_{4},\bar{e}_{4}\wedge\bar{e}_{2}\}$
of $\wedge^{2}T\Sigma^{3}$, where $\{\bar{e}_{2},\bar{e}_{3},\bar{e}_{4}\}$ is a local orthonormal frame of $T\Sigma^{3}$. Suppose that, in this frame, $\overline{Rm}$ is diagonal with entries
$\bar{R}_{2323}=:m_{1},\quad\bar{R}_{2424}=:m_{2},\quad\bar{R}_{3434}=:m_{3}$
such that $m_{1}\leq m_{2}\leq m_{3}$. Then, with respect to the tangent frame $\{\bar{e}_{2},\bar{e}_{3},\bar{e}_{4}\}$, the Ricci tensor $\overline{R}c$} of $\Sigma^{3}$ takes the form
$\overline{Rc}=\begin{pmatrix}m_{1}+m_{2}&0&0\\ 0&m_{1}+m_{3}&0\\ 0&0&m_{2}+m_{3}\end{pmatrix}$.
The scalar curvature of $\Sigma^{3}$ is given by
$\bar{R}=2(m_{1}+m_{2}+m_{3}).$
For the cone $\mathcal{C}^{4}$, we choose
$e_{1}=\partial_{r},\qquad e_{i}=r^{-1}\bar{e}_{i}\ (i=2,3,4),$
so that $\{e_{1},e_{2},e_{3},e_{4}\}$ forms an orthonormal frame of $T\mathcal{C}$ with respect to $g_{c}$. Then, by (6.1), we have
$R_{1jkl}$ $=0,\quad 1\leq j,k,l\leq 4,$
$R_{abcd}$ $=r^{-2}\left[\bar{R}_{abcd}-\big(\bar{g}_{ac}\bar{g}_{bd}-\bar{g}_{ad}\bar{g}_{bc}\big)\right],\quad 2\leq a,b,c,d\leq 4$
where $\bar{R}_{abcd}$ denotes the curvature tensor of the link $(\Sigma^{3},\bar{g})$.
Therefore, by (6.6), we obtain
$A_{11}=C_{11}$ $=\tfrac{1}{2}R_{3434}=\tfrac{1}{2r^{2}}(m_{3}-1),$ (6.8)
$A_{22}=C_{22}$ $=\tfrac{1}{2}R_{2424}=\tfrac{1}{2r^{2}}(m_{2}-1),$
$A_{33}=C_{33}$ $=\tfrac{1}{2}R_{2323}=\tfrac{1}{2r^{2}}(m_{1}-1),$
and $A_{ij}=C_{ij}=0\ (i\neq j)$, e.g.,
$A_{12}=C_{12}=\tfrac{1}{2}R_{3442}=0,\quad\text{etc.}$
Similarly, by (6.7), we have
$B_{11}$ $=-\tfrac{1}{2}R_{3434}=-\tfrac{1}{2r^{2}}(m_{3}-1)=-A_{11},$ (6.9)
$B_{22}$ $=-\tfrac{1}{2}R_{4242}=-\tfrac{1}{2r^{2}}(m_{2}-1)=-A_{22},$
$B_{33}$ $=-\tfrac{1}{2}R_{2323}=-\tfrac{1}{2r^{2}}(m_{1}-1)=-A_{33},$
and $B_{ij}=0\ (i\neq j)$, e.g.,
$B_{12}=-\tfrac{1}{2}R_{3442}=0,\quad\text{etc.}$
Moreover, for the Ricci tensor of $(\mathcal{C},g_{c})$, we have $R_{ij}=0\ (i\neq j)$ and
$R_{11}$ $=0,$ $R_{22}=r^{-2}\left(m_{1}+m_{2}-2\right),$
$R_{33}$ $=r^{-2}\left(m_{1}+m_{3}-2\right),$ $R_{44}=r^{-2}\left(m_{2}+m_{3}-2\right).$
In conclusion, based on the above computations, we have
Lemma 6.3.
Let $\mathcal{C}^{4}:=[0,\infty)\times\Sigma^{3}$ be a $4$-dimensional cone equipped with the Riemannian metric $g_{c}=dr^{2}+r^{2}\bar{g}_{\Sigma}$. Then, the following statements hold:
  • (i)
    $(\mathcal{C}^{4},g_{c})$ has $A\geq 0$ if and only it has nonnegative curvature operator $Rm\geq 0$, if and only if $\overline{Rm}\geq Id$.
  • (ii)
    $(\mathcal{C}^{4},g_{c})$ has half WPIC if and only if it has WPIC and nonnegative Ricci curvature $Rc\geq 0$, if and only if $\overline{Rc}\geq 2{\bar{g}}$.

### v3

@@
 Abstract.
-Since the well-known work of Hamilton [62] and Ivey [64], the Hamilton-Ivey curvature pinching and its generalizations have become a signature feature of gradient shrinking and steady Ricci solitons, and more generally, of ancient solutions to the Ricci flow.
-However, analogous results for gradient expanding Ricci solitons have remained elusive.
-This is largely due to the fact that the proofs of existing curvature pinching estimates crucially rely on shrinking and steady solitons being ancient, a property not shared by gradient Ricci expanders.
-In this paper, we investigate curvature pinching phenomena in non-compact asymptotically conical gradient expanding Ricci solitons and establish several Hamilton-Ivey type curvature pinching estimates.
+In this paper, we investigate curvature pinching phenomena in complete non-compact asymptotically conical gradient expanding Ricci solitons and establish several Hamilton–Ivey type curvature pinching estimates.
 These results are parallel to those known for shrinking and steady Ricci solitons.
-In particular, we prove a three-dimensional Hamilton-Ivey type curvature pinching theorem: any three-dimensional non-compact gradient Ricci expander, which is asymptotic to a cone with positive scalar curvature, must have positive sectional curvature.
-As an application, we combine our result with that of Deruelle [51] to establish a uniqueness theorem for three-dimensional asymptotically conical expanders with positive scalar curvature.
-Furthermore, we formulate a general proof method and apply it to obtain analogues of several additional known generalized Hamilton-Ivey type curvature pinching results for ancient solutions.
+In particular, we prove a three-dimensional Hamilton–Ivey type curvature pinching theorem: any three-dimensional non-compact gradient Ricci expander, which is asymptotic to a cone with positive scalar curvature, must have positive sectional curvature.
+Furthermore, we formulate a general method and apply it to obtain analogues of several additional known generalized Hamilton–Ivey type curvature pinching results for ancient solutions.
 Among these is a curvature pinching estimate for four-dimensional asymptotically conical Ricci expanders with uniformly positive isotropic curvature, analogous to a result for four-dimensional gradient steady solitons due to Brendle [8].
@@
 In particular, shrinking and steady solitons often arise as Type I and, respectively, Type II singularity models [63, 77, 56, 17] in the Ricci flow, while expanding solitons may arise as Type III singularity models [17, 38] and over which the matrix Li-Yau-Hamilton (LYH) differential Harnack inequality [62, 15] becomes equality.
-The first examples of gradient expanding Ricci solitons are the one-parameter family of complete, rotationally symmetric, asymptotically conical gradient expanding Ricci solitons on $\mathbb{R}^{n}$ ($n\geq 3$) with positive or negative sectional curvature constructed by Bryant [11], and the one-parameter family of complete, $U(n)$-invariant, asymptotically conical gradient expanding Kähler-Ricci solitons on $\mathbb{C}^{n}$ with similar curvature behavior constructed by the first author [16].
+The first examples of gradient expanding Ricci solitons are the one-parameter family of complete, rotationally symmetric, asymptotically conical gradient expanding Ricci solitons on $\mathbb{R}^{n}$ ($n\geq 3$) with positive (and negative) sectional curvature constructed by Bryant [12], and the one-parameter family of complete, $U(n)$-invariant, asymptotically conical gradient expanding Kähler-Ricci solitons on $\mathbb{C}^{n}$ with similar curvature behavior constructed by the first author [17].
 The constructions in [16, 17] were later extended by Feldman-Ilmanen-Knopf [57] to produce gradient expanding Kähler-Ricci solitons on the complex line bundles O(k)O(-k) ($k>n$) over the complex projective space ${\mathbb{C}}P^{n}$ ($n\geq 1$), and further generalized by Dancer-Wang [51].
-Additional constructions can be found in [65, 2, 49, 57, 12, 85, 86, 77, 41, 42].
+Additional constructions can be found in [66, 2, 50, 58, 13, 86, 87, 78, 43].
 Asymptotically conical gradient expanding Ricci solitons have received increasing attention in recent years.
@@
 Curvature estimates for gradient expanding Ricci solitons with $Rc\geq 0$ or scalar curvature $R>0$ have also been established in [24, 25], mirroring those for shrinking solitons [75, 76] or steady solitons [23, 32, 18].
-Despite the progress described above, a key feature well known for gradient shrinking and steady Ricci solitons has remained absent in the expanding case: Hamilton-Ivey type curvature pinching.
-As is well-known, a hallmark of the three-dimensional Ricci flow is the Hamilton-Ivey curvature pinching [62, 64] (see also [27, Theorem 2.4.1]), which asserts that when curvature blows up, the positive part blows up at a faster rate than the absolute value of the negative part.
+Despite the progress described above, a key feature well known for gradient shrinking and steady Ricci solitons has remained absent in the expanding case: Hamilton–Ivey type curvature pinching.
+As is well-known, a hallmark of the three-dimensional Ricci flow is the Hamilton–Ivey curvature pinching [63, 65] (see also [29, Theorem 2.4.1]), which asserts that when curvature blows up, the positive part blows up at a faster rate than the absolute value of the negative part.
 In particular, 3-dimensional singularity models that are shrinking or steady gradient Ricci solitons, or more generally ancient solutions, must have nonnegative sectional curvature.
-This remarkable feature is powerful: it enables the application of the Li-Yau-Hamilton differential Harnack inequality and special geometry of nonnegatively curved 3-manifolds to effectively analyze three-dimensional singularity models.
-The Hamilton-Ivey curvature pinching was later extended by B.-L.
+This feature is remarkable: it enables the application of the powerful Li–Yau–Hamilton differential Harnack inequality and the geometry of nonnegatively curved 3-manifolds to effectively analyze three-dimensional singularity models.
+The Hamilton–Ivey curvature pinching was later extended by B.-L.
 Chen [37] to arbitrary 3-dimensional ancient solutions [^1] Ancient solutions exist for all past time, up to a final singular time; immortal solutions, starting at some initial time, exist for all future time., which include gradient shrinking and steady solitons as important special cases.
@@
 This has played a crucial role in the classifications of 3-dimensional gradient shrinking and steady Ricci solitons [22, 7, 67, 68, 69], as well as 3D ancient solutions [9, 10].
-In higher dimensions, various forms of the generalized Hamilton-Ivey curvature pinching have been established for ancient solutions under suitable assumptions, see, e.g., [89, 3, 70, 43].
-More recently, the authors proved certain curvature pinching properties for $4$-dimensional ancient solutions with positive isotropic curvature (PIC) or half PIC [24, 25], leading to partial classifications of $4$-dimensional shrinking and steady gradient Ricci solitons with weakly positive isotropic curvature (WPIC) or half WPIC.
+In higher dimensions, various forms of the generalized Hamilton–Ivey curvature pinching have been established for ancient solutions under suitable assumptions, see, e.g., [90, 3, 71, 44].
+More recently, the authors proved certain curvature pinching properties for $4$-dimensional ancient solutions with positive isotropic curvature (PIC) or half-PIC [26, 27], leading to partial classifications of $4$-dimensional shrinking and steady gradient Ricci solitons with weakly positive isotropic curvature (WPIC) or half-WPIC.
 However, analogous curvature pinching results for gradient expanding Ricci solitons have remained largely unknown.
-This is partly because the proofs of Hamilton-Ivey-type estimates for shrinking and steady solitons rely in an essential way on the ancient nature of these solutions.
+This is partly because the proofs of Hamilton–Ivey-type estimates for shrinking and steady solitons rely in an essential way on the ancient nature of these solutions.
 Since gradient expanders are not ancient solutions but special immortal solutions, those methods do not apply.
-Motivated in part by the second author’s work [87] on the convexity of mean convex, asymptotically conical self-expanders in mean curvature flow, and in part by the close resemblance between curvature estimates for Ricci expanders with nonnegative curvature [22, 23] and those for shrinking or steady Ricci solitons [74, 75, 21, 30, 17] (see especially the comparison in dimension four given in [17]), we began to investigate whether analogous Hamilton-Ivey-type curvature pinching properties might hold in the expanding case.
-By adapting an argument from the second author’s recent work [87] on asymptotically conical mean curvature expanders – inspired in turn by Spruck-Xiao [84] and Xie-Yu [88] – we have indeed established a number of generalized Hamilton-Ivey-type curvature pinching results for non-compact, asymptotically conical, gradient expanding Ricci solitons with positive scalar curvature.
-Our first result is a Hamilton-Ivey type curvature pinching for $3$-dimensional non-compact, asymptotically conical, gradient expanders with positive scalar curvature.
+Motivated in part by the second author’s work [88] on the convexity of mean convex, asymptotically conical self-expanders in mean curvature flow, and in part by the close resemblance between curvature estimates for Ricci expanders with nonnegative curvature [24, 25] and those for shrinking or steady Ricci solitons [75, 76, 23, 32, 18] (see especially the comparison in dimension four given in [18]), we began to investigate whether analogous Hamilton–Ivey-type curvature pinching properties might hold in the expanding case.
+By adapting an argument from the second author’s recent work [88] on asymptotically conical mean curvature expanders–inspired in turn by Spruck-Xiao [85] and Xie-Yu [89]–we have indeed established a number of generalized Hamilton–Ivey-type curvature pinching results for non-compact, asymptotically conical, gradient expanding Ricci solitons with positive scalar curvature.
+Our first result is a Hamilton–Ivey type curvature pinching for $3$-dimensional complete non-compact, asymptotically conical, gradient Ricci expanders with positive scalar curvature.
 Theorem 1.1.
-Let $(M^{3},g,f)$ be a $3$-dimensional non-compact, asymptotically conical gradient expanding Ricci soliton.
+Let $(M^{3},g,f)$ be a $3$-dimensional non-compact, asymptotically conical [^2] For all our results in this paper, except Corollary 1.2, it suffices to assume $C^{2}$-asymptotics in the sense of Definition 2.1(a).
+gradient expanding Ricci soliton.
 Suppose the asymptotic cone has positive scalar curvature.
@@
 Chen’s result for three-dimensional ancient solutions to the Ricci flow [37].
-Moreover, after our paper was completed, the authors learned from P.-Y.
-Chan that a similar result to Theorem 1.1 was proved in [34, Corollary 1.16] by a different method.
-As an application of Theorem 1.1, for $n=3$, we can extend Deruelle’s uniqueness result [51, Theorem 1.3] to asymptotically conical expanding Ricci solitons with positive scalar curvature.
-Corollary 1.1.
-Let $(\Sigma^{2},\bar{g})$ be a smooth 2-sphere with Gaussian curvature $\bar{K}>1$.
-Then, there exists a unique $3$-dimensional gradient expanding Ricci soliton with positive scalar curvature asymptotic to the cone $\mathcal{C}^{3}:=[0,\infty)\times\Sigma^{2}$ with link $\Sigma^{2}$.
-For $n\geq 4$, we have the following Hamilton-Ivey type curvature pinching result for asymptotically conical, gradient expanding solitons with positive scalar curvature and vanishing Weyl tensor.
+Moreover, after our paper was completed, we learned from P.-Y.
+Chan that a similar result to Theorem 1.1 was proved in [36, Corollary 1.16] by a different method.
+For $n\geq 4$, we have the following Hamilton–Ivey type curvature pinching result for asymptotically conical, locally conformally flat gradient expanding solitons with positive scalar curvature.
 Theorem 1.2.
@@
 Zhang’s result for locally conformally flat gradient shrinking (and steady) solitons [90].
-Moreover, by the work of [19, 26], the locally conformally flat assumption can be replaced by the weaker assumption of vanishing $D$-tensor introduced in [18, 19].
-As a consequence, by combining with the work of Cao-Chen [19] and Chen-Wang [39], we have the following application in dimension four.
+Moreover, by [21, 28], the locally conformally flat assumption can be replaced by the weaker assumption of the vanishing $D$-tensor introduced in [20, 21].
+As a consequence, by combining Theorem 1.2 with the work of Cao-Chen [21] and Chen-Wang [41], we have the following application in dimension four.
 Corollary 1.1.
-Let $(M^{4},g,f)$ be a $4$-dimensional non-compact, half conformally flat, asymptotically conical gradient expanding Ricci soliton.
+Let $(M^{4},g,f)$ be a $4$-dimensional non-compact, half-conformally flat, asymptotically conical gradient expanding Ricci soliton.
 Suppose the asymptotic cone has positive scalar curvature.
 Then, $(M^{4},g,f)$ has positive curvature operator.
-Our next two results concern curvature pinching of 4-dimensional asymptotically conical gradient expanding Ricci solitons with either positive isotropic curvature (PIC) or half positive isotropic curvature (half-PIC); see Section 2.2 for definitions.
+Remark 1.3.
+By [19, Theorem 5.8], the expanding solitons in Theorem 1.2 and Corollary 1.1 are rotationally symmetric; see also Cao-Yu [28, Corollary 3.4].
+Our next two results concern curvature pinching of 4-dimensional asymptotically conical gradient expanding Ricci solitons with either positive isotropic curvature (PIC) or half-positive isotropic curvature (half-PIC).
+Recall that, for any oriented 4-manifold $(M^{4},g)$, $2$-forms admit the decomposition $\wedge^{2}(M)=\wedge^{+}(M)\oplus\wedge^{-}(M)$, into self-dual and anti-self-dual $2$-forms.
+Accordingly, the Riemann curvature operator admits a block decomposition
+$Rm=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\frac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\frac{R}{12}I\end{pmatrix}$,
+where $W^{\pm}$ denote the self-dual and anti-self-dual part of the Weyl tensor, and $\mathring{Rc}$ the traceless Ricci part.
+It turns out that $(M^{4},g)$ has PIC (half-PIC) if and only if $A$ and $C$ (AA or $C$) are 2-positive [64]; see Section 2.2 for more details.
 Theorem 1.3.
-Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with half positive isotropic curvature.
-If the asymptotic cone has positive scalar curvature and satisfies either $A\geq 0$ or $C\geq 0$, then $(M^{4},g,f)$ has either $A>0$ or $C>0$.
+Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with half-positive isotropic curvature (half-weakly PIC).
+If the asymptotic cone has positive scalar curvature and satisfies either $A\geq 0$ or $C\geq 0$, then $(M^{4},g,f)$ has $A>0$ ($A\geq 0$) or $C>0$ ($C\geq 0$).
 Theorem 1.4.
-Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with positive isotropic curvature.
-If the asymptotic cone has positive scalar curvature, then the Ricci curvature of $(M^{4},g,f)$ is $2$-positive.
+Let $(M^{4},g,f)$ be a $4$-dimensional non-compact asymptotically conical gradient expanding Ricci soliton with positive isotropic curvature (weakly PIC).
+If the asymptotic cone has positive scalar curvature, then the Ricci curvature of $(M^{4},g,f)$ is $2$-positive (22-nonnegative), and $|Rm|\leq 2R$.
 Remark 1.4.
-Theorems 1.3 and 1.4 form an analogue of the authors’ previous results for $4$-dimensional complete ancient solutions with half PIC and PIC [24, 25], respectively.
+Theorems 1.3 and 1.4 are analogues of our previous results for $4$-dimensional complete ancient solutions with half-PIC [26, Proposition 3.1] and PIC [27, Theorem 1.1], respectively.
 Moreover, both results are valid under some slightly weaker assumptions; see Theorem 4.1 and Theorem 4.2 in Section 4.
-By observing common patterns in the proofs of Theorem 1.3 and Theorem 1.4, we formulate a general method of proof (Lemma 5.1) and apply it to obtain analogues of several additional known generalized Hamilton-Ivey type curvature pinching results for ancient solutions.
-This includes the following result for $4$-dimensional asymptotically conical gradient expanding Ricci soliton with uniformly positive isotropic curvature (UPIC), as well as several others stated in Theorem 5.1.
+By observing the common pattern in the proofs of Theorem 1.3 and Theorem 1.4, we formulate a general method of proof (Lemma 5.1) and apply it to obtain analogues of several additional known generalized Hamilton–Ivey type curvature pinching results for ancient solutions.
+This includes the following result for $4$-dimensional asymptotically conical gradient expanding Ricci solitons with uniformly positive isotropic curvature (UPIC), as well as several other results as stated in Theorem 5.1.
 Theorem 1.5.
@@
 Let $(M^{4},g,f)$ be a $4$-dimensional non-compact gradient expanding Ricci soliton with uniformly positive isotropic curvature.
-If $(M^{4},g,f)$ is asymptotic to a non-flat Euclidean cone, then it is rotationally symmetric.
+If $(M^{4},g,f)$ is asymptotic [^3] Here, the notion of asymptotically conical is $C^{2}$-asymptotic in the sense of Chodosh [45, Definition 1.1] which is slightly stronger than that in our Definition 2.1.
+to a non-flat Euclidean cone, then it is rotationally symmetric.
+Remark 1.6.
+We note that any $4$-dimensional non-compact gradient expanding Ricci soliton with PIC that is asymptotic to a non-flat Euclidean cone automatically has UPIC.
+Therefore, one could replace the assumption of UPIC by PIC in both Theorem 1.5 and Corollary 1.2.
 Organization of the Paper.
@@
 Chan for very helpful information that allowed the removal of the positive scalar curvature assumption on $(M^{n},f,g)$ in Theorem 1.1 and Theorem 1.2 in an earlier version of the paper.
+We also thank the anonymous referees for their helpful comments.
 2. Preliminaries
@@
 where $\wedge^{+}(M)$ and $\wedge^{-}(M)$ denote the subbundles of self-dual and anti-self-dual $2$-forms, respectively. With respect to this splitting, the curvature operator has block form
-${Rm}=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\tfrac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\tfrac{R}{12}I\end{pmatrix}$,
+${Rm}=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\tfrac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\tfrac{R}{12}I\end{pmatrix}$, (2.1)
 […]
-Let $A_{1}\leq A_{2}\leq A_{3}$ and $C_{1}\leq C_{2}\leq C_{3}$ denote the eigenvalues of $A$ and $C$, respectively, and let $a_{1}\leq a_{2}\leq a_{3}$ and $c_{1}\leq c_{2}\leq c_{3}$ denote the eigenvalues of $W^{+}$ and $W^{-}$, respectively.
-Then
-$A_{i}=a_{i}+\tfrac{R}{12},\qquad C_{i}=c_{i}+\tfrac{R}{12},\qquad(i=1,2,3).$
-Since $a_{1}+a_{2}+a_{3}=\operatorname{tr}W^{+}=0=\operatorname{tr}W^{-}=c_{1}+c_{2}+c_{3}$, it follows that
-$\operatorname{tr}A=\operatorname{tr}C=\tfrac{R}{4}.$
+Let $A_{1}\leq A_{2}\leq A_{3}$ and $C_{1}\leq C_{2}\leq C_{3}$ denote the eigenvalues of $A$ and $C$, respectively. It is also well-known that $\operatorname{tr}A=\operatorname{tr}C={R}/{4}$.
 Definition 2.2.
@@
 It also plays a key role in the convergence theory for the higher dimensional Ricci flow, especially in Brendle-Schoen’s proof of the 1/4-pinching differentiable sphere theorem [11].
-It turns out that, in dimension four, these curvature conditions (and their natural extensions, half PIC or half WPIC) can be characterized in terms of the $3\times 3$ matrices $A$ and $C$ as follows:
+It turns out that, in dimension four, these curvature conditions (and their natural extensions, half-PIC or half-WPIC) can be characterized in terms of the $3\times 3$ matrices $A$ and $C$ as follows:
   • PIC (WPIC) if and only if $A$ and $C$ are 2-positive (weakly 2-positive), i.e., $A_{1}+A_{2}>0$ ($A_{1}+A_{2}\geq 0$) and $C_{1}+C_{2}>0$ ($C_{1}+C_{2}\geq 0$) on $M$ [62];
-  • half PIC (half WPIC) if and only if either $A$ or $C$ is 2-positive (weakly 2-positive), i.e., $A_{1}+A_{2}>0$ ($A_{1}+A_{2}\geq 0$) or $C_{1}+C_{2}>0$ ($C_{1}+C_{2}\geq 0$).
+  • half-PIC (half-WPIC) if and only if either $A$ or $C$ is 2-positive (weakly 2-positive), i.e., $A_{1}+A_{2}>0$ ($A_{1}+A_{2}\geq 0$) or $C_{1}+C_{2}>0$ ($C_{1}+C_{2}\geq 0$).
 Definition 2.3.
@@
 $\Delta_{f}Rm=2\rho Rm-2(Rm^{2}+Rm^{\sharp}),$
-$\Delta_{f}A=2\rho A-2(A^{2}+2A^{\sharp}+B{}^{t}B),$
+$\Delta_{f}A=2\rho A-2(A^{2}+2A^{\sharp}+BB^{t}),$
 $\Delta_{f}B=2\rho B-2(AB+BC+2B^{\sharp}),$
-$\Delta_{f}C=2\rho C-2(C^{2}+2C^{\sharp}+{}^{t}BB).$
+$\Delta_{f}C=2\rho C-2(C^{2}+2C^{\sharp}+B^{t}B).$
 Here, for any $3\times 3$ matrix $D$, we denote by $D^{2}$ its square and by $D^{\sharp}$ the transpose of its adjoint.
@@
 Remark 2.1.
-Except for the first identity in Lemma 2.1, the factor 22 differs from [60] due to our normalization of the inner product on $\wedge^{2}(M)$ (see [24, (2.4)]).
+Except for the first identity in Lemma 2.1, the factor 22 differs from [61] due to our normalization of the inner product on $\wedge^{2}(M)$ (see (A.6)).
 Moreover, the first two equations are valid in all dimensions.
@@
 3. Curvature pinching for asymptotically conical Ricci expanders with vanishing Weyl tensor
-This section is devoted to the proof of Theorem 1.1 and Theorem 1.2. For the reader’s convenience, we restate Theorem 1.2 in the following form.
+This section is devoted to the proof of Theorem 1.1 and Theorem 1.2. Note that both theorems follow from the following result.
 Theorem 3.1.
@@
 Then, $(M^{n},g,f)$ has positive curvature operator.
-First of all, since the asymptotic cone of $(M^{n},g,f)$ has positive scalar curvature, it follows from the strong maximum principle and [31, Theorem 1.6] that $(M^{n},g,f)$ itself has positive scalar curvature $R>0$.
+First of all, since the asymptotic cone of $(M^{n},g,f)$ has positive scalar curvature, it follows from the strong maximum principle and [33, Theorem 1.6] that $(M^{n},g,f)$ itself has positive scalar curvature $R>0$.
 This fact will be used in the proofs of Theorem 3.1 and Theorem 1.1.
-Next, shall need the following differential inequality for the ratio between the smallest Ricci eigenvalue and the scalar curvature.
+Next, we shall need the following differential inequality for the ratio between the smallest Ricci eigenvalue $\lambda_{1}$ and the scalar curvature $R$.
 Lemma 3.1 ([55, 81]).
@@
 Proof.
-Fix any point $p_{0}\in M$, and choose a unit vector field $e_{1}$ such that $Rc(e_{1})=\lambda_{1}e_{1}$ at $p_{0}$, where $\lambda_{1}$ denotes the smallest Ricci eigenvalue.
+Let $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$ be the Ricci eigenvalues.
+Fix any point $p_{0}\in M$, and choose a unit vector field $e_{1}$ such that $Rc(e_{1})=\lambda_{1}e_{1}$ at $p_{0}$.
 Extend $e_{1}$ by parallel transport along geodesics emanating from $p_{0}$.
@@
 $\begin{split}\Delta_{f}\lambda_{1}&\leq\Delta_{f}Rc(e_{1},e_{1})\\ &=(\Delta_{f}Rc)(e_{1},e_{1})\\ &=2\rho\lambda_{1}-\frac{2nR}{(n-1)(n-2)}\lambda_{1}+\frac{4}{n-2}\lambda_{1}^{2}-\frac{2}{n-2}\left(|Rc|^{2}-\frac{R^{2}}{n-1}\right),\end{split}$ (3.1)
-where in the last equality, we have used [54, (2.8)] (see also [80, Lemma 2.5]).
+where in the last equality, we have used [55, (2.8)] (see also [81, Lemma 2.5]).
 On the other hand, we have
@@
 $\begin{split}h_{1}&=S^{2}\big[S+(3-n)\lambda_{1}\big]+(2-n)\lambda_{1}^{2}\big[S-(n-1)\lambda_{1}\big]\\ &\quad+(n-1)T\big((n-3)\lambda_{1}-S\big)\\ &=(n-2)\lambda_{1}^{2}(n\lambda_{1}-R)+\big[(n-2)\lambda_{1}-R\big]\big[(n-1)T-S^{2}\big].\end{split}$
-Finally, it is clear from the above expression that $h\leq 0$. ∎
-Now, we divide the proof of Theorem 3.1 into two parts.
-Part I: $(M^{n},g,f)$ has nonnegativity Ricci curvature $Rc\geq 0$.
+Finally, it is clear from the above expression that $h_{1}\leq 0$. ∎
+Now, we divide the proof of Theorem 3.1 into three parts.
+Part I: $(M^{n},g,f)$ has nonnegative Ricci curvature $Rc\geq 0$.
 Proof of $Rc\geq 0$.
@@
 By Lemma 3.1 and Calabi’s barrier strong maximum principle (Lemma 2.2), the ratio $\lambda_{1}/R$ must be constant on some bounded connected domain $p_{0}\in\Omega\subset M^{-}$.
-In particular, $h\equiv 0$ on $\Omega$ which forces
+In particular, $h_{1}\equiv 0$ on $\Omega$ which forces
 $\lambda_{1}^{2}(n\lambda_{1}-R)\equiv 0,\qquad\bigl((n-2)\lambda_{1}-R\bigr)\Biggl((n-1)\sum_{j=2}^{n}\lambda_{j}^{2}-\Bigl(\sum_{j=2}^{n}\lambda_{j}\Bigr)^{2}\Biggr)\equiv 0.$ (3.4)
@@
 On the other hand, since the expanding Ricci soliton satisfies $W\equiv 0$, its asymptotic cone $\mathcal{C}$ also has vanishing Weyl tensor.
-By Lemma 6.2, $\mathcal{C}$ has nonnegative Ricci curvature, whose smallest eigenvalue $\tilde{\lambda}_{1}\equiv 0$ occurs in the radial direction.
+By Lemma A.1 (for $n=3$) and Lemma A.2 (for $n\geq 4$), $\mathcal{C}$ has nonnegative Ricci curvature, whose smallest eigenvalue $\tilde{\lambda}_{1}\equiv 0$ occurs in the radial direction.
 Hence
@@
 Combining both cases, we conclude that $M^{-}=\emptyset$. Therefore, $(M^{n},g,f)$ has nonnegative Ricci curvature. ∎
-Part II: $(M^{n},g,f)$ has positive curvature operator $Rm>0$.
-First of all, when the Weyl tensor $W$ vanishes, we may express the smallest sectional curvature as (see, e.g., [89, Proposition 3.1])
+Part II: $(M^{n},g,f)$ has nonnegative curvature operator $Rm\geq 0$.
+First of all, when the Weyl tensor $W$ vanishes, we may express the smallest sectional curvature as (see, e.g., [90, Proposition 3.1])
 $\mu:=\min_{i,j}R_{ijij}=R_{1212}=\frac{1}{n-2}\big[\lambda_{1}+\lambda_{2}-\frac{R}{n-1}\big],$
@@
 Remark 3.1.
-We also note that, for any manifold with vanishing Weyl tensor, $\mu$ also coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [89, Proposition 3.1(i)]).
+We also note that, for any manifold with vanishing Weyl tensor, $\mu$ also coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [90, Proposition 3.1(i)]).
 In particular, when $W=0$, nonnegative sectional curvature is equivalent to nonnegative curvature operator $Rm\geq 0$.
@@
 with equality at $p_{0}$, i.e., $\tilde{\lambda}$ is a barrier function for $\lambda$.
-Evaluating at $p_{0}$ in the barrier sense, using (3.6) and [54, (2.8)] (see also [80, Lemma 2.5]), we obtain
+Evaluating at $p_{0}$ in the barrier sense, using (3.6) and [55, (2.8)] (see also [81, Lemma 2.5]), we obtain
 $\begin{split}\Delta_{f}\lambda&=\Delta_{f}(\lambda_{1}+\lambda_{2})\\ &\leq\Delta_{f}\big(Rc(\tilde{e}_{1},\tilde{e}_{1})+Rc(\tilde{e}_{2},\tilde{e}_{2})\big)\\ &=(\Delta_{f}Rc)(\tilde{e}_{1},\tilde{e}_{1})+(\Delta_{f}Rc)(\tilde{e}_{2},\tilde{e}_{2})\\ &=2\rho\lambda_{1}-\frac{2nR}{(n-1)(n-2)}\lambda_{1}+\frac{4}{n-2}\lambda_{1}^{2}-\frac{2}{n-2}\left(|Rc|^{2}-\frac{R^{2}}{n-1}\right)\\ &\quad+2\rho\lambda_{2}-\frac{2nR}{(n-1)(n-2)}\lambda_{2}+\frac{4}{n-2}\lambda_{2}^{2}-\frac{2}{n-2}\left(|Rc|^{2}-\frac{R^{2}}{n-1}\right).\end{split}$
@@
 One can notice that every term, except possibly the first one and the last one, in the above expression (3.7) of $E$ is nonpositive. On the other hand, the last term can be rewritten as
-$\begin{split}\bigl((n-2)\lambda_{2}-R\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}&=-\Bigg(\lambda_{1}+\lambda_{2}+\sum_{j=3}^{n}(\lambda_{j}-\lambda_{2})\Bigg)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &=-(\lambda_{1}+\lambda_{2})\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &\quad-\Biggl(\sum_{k=3}^{n}(\lambda_{k}-\lambda_{2})\Biggr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\end{split}$ (3.8)
+$\begin{split}\bigl((n-2)\lambda_{2}-R\bigr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}&=-\Bigg(\lambda_{1}+\lambda_{2}+\sum_{j=3}^{n}(\lambda_{j}-\lambda_{2})\Bigg)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &=-(\lambda_{1}+\lambda_{2})\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}\\ &\quad-\Biggl(\sum_{k=3}^{n}(\lambda_{k}-\lambda_{2})\Biggr)\sum_{\begin{subarray}{c}i>j\\ i,j\neq 2\end{subarray}}(\lambda_{i}-\lambda_{j})^{2}.\end{split}$ (3.8)
 Combining (3.7) and (3.8), we arrive at the desired expression of $E$.
@@
 ∎
-Now, we are ready to complete the proof of Theorem 3.1, i.e., $(M^{n},g,f)$ has positive curvature operator $Rm>0$.
-Proof of $Rm>0$.
+Now, we are ready to complete the proof that $(M^{n},g,f)$ has $Rm\geq 0$.
+Proof of $Rm\geq 0$.
 First, we claim that $(M^{n},g,f)$ must have nonnegative sectional curvature. Suppose not. Then, as before, we argue by contradiction. Assume the claim fails. Then the set
@@
 Suppose the negative infimum $\epsilon_{0}$ is attained at some point $p_{0}\in M$.
-By Theorem 3.1 (i), we know that the Ricci curvature is nonnegative.
+By Part I of the proof above, we know that the Ricci curvature is nonnegative.
 Then, by Lemma 3.2, we have
@@
 $\bullet$ Subcase 1a: $\lambda_{2}=0$ at $p_{0}$.
-By Theorem 3.1 (i), we have $\operatorname{Ric}\geq 0$, which implies $\lambda_{1}(p_{0})=0$.
-Then, the third relation in (3.10) yields $\lambda_{i}=\lambda_{j}$ at $p_{0}$ for all $i,j\geq 2$, hence $\lambda_{i}(p_{0})=0$ for all $i$, contradicting the assumption that the scalar curvature $R$ is positive.
+By Part I, we have $\operatorname{Ric}\geq 0$, which implies $\lambda_{1}(p_{0})=0$.
+Then, the third relation in (3.10) yields $\lambda_{i}=\lambda_{j}$ at $p_{0}$ for all $i,j\geq 2$, hence $\lambda_{i}(p_{0})=0$ for all $i$, contradicting the fact that the scalar curvature $R$ is positive.
 $\bullet$ Subcase 1b: $\lambda_{2}=\lambda_{j}$, at $p_{0}$, for all $j\geq 3$.
@@
 Suppose $\mu/R$ attains its negative infimum at infinity.
-Then, by Lemma 6.2 and essentially the same argument as in the proof of $Rc\geq 0$ in Part I, we can rule out Case 2.
+Then, by Lemma A.1 and Lemma A.2 and essentially the same argument as in the proof of $Rc\geq 0$ in Part I, we can rule out Case 2.
 We omit the details.
@@
 As nonnegative sectional curvature is equivalent to nonnegative curvature operator for manifolds with vanishing Weyl tensor $W=0$, it follows that the expanding soliton $(M,g,f)$ has nonnegative curvature operator $Rm\geq 0$.
-In fact, $\mu$ coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [89, Proposition 3.1(i)]).
-Finally, it remains to prove $(M,g,f)$ has positive curvature operator $Rm>0$.
-To begin with, following the proof of [25, Theorem 1.3], we observe that the null space of the curvature operator, $\ker(Rm)$, is invariant under parallel translation.
+In fact, $\mu$ coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [90, Proposition 3.1(i)]).
+∎
+Part III: $(M^{n},g,f)$ has positive curvature operator $Rm>0$.
+Proof of $Rm>0$.
+To begin with, following the proof of [27, Theorem 1.3], we observe that the null space of the curvature operator, $\ker(Rm)$, is invariant under parallel translation.
 Indeed, consider the canonical immortal solution
@@
 and the fact that $(M,g)$ has nonnegative curvature operator $Rm\geq 0$, it follows that the quadratic term $Rm(t)^{2}+Rm(t)^{\sharp}\geq 0$ for all $t\in[0,1]$.
-Thus, by Hamilton’s strong maximum principle (see, e.g., [27, Theorem 2.2.1]), there exists an interval $0<t<\delta$ over which the rank of $Rm(t)$ is constant, and $\ker(Rm(t))$ is invariant under parallel translation.
+Thus, by Hamilton’s strong maximum principle (see, e.g., [29, Theorem 2.2.1]), there exists an interval $0<t<\delta$ over which the rank of $Rm(t)$ is constant, and $\ker(Rm(t))$ is invariant under parallel translation.
 Since $g(t)=(1+t)\Phi(t)^{\ast}(g)$ with $g(0)=g$, we conclude that the rank of $Rm=Rm(0)$ is locally constant and that $\ker(Rm)=\ker(Rm(0))$ is invariant under parallel translation.
@@
 Moreover, as $\ker(Rm)$ is invariant under parallel translation, there exists an orthogonal decomposition of the tangent bundle $TU=V_{1}\oplus V_{2}$, where $V_{1}=\ker(Rm)$ with $\dim(V_{1})\geq 1$, and both $V_{1}$ and $V_{2}$ are invariant under parallel translation.
-By [60, Lemma 9.1], $M$ locally splits off at least one Euclidean factor.
+By [61, Lemma 9.1], $M$ locally splits off at least one Euclidean factor.
 Since expanding Ricci solitons are real analytic, completeness implies that this local splitting extends globally.
@@
 Proof of Corollary 1.1.
-Let $(M^{4},g,f)$ be a non-compact, half conformally flat (i.e., either $W^{+}=0$ or $W^{-}=0$), asymptotically conical Ricci expander with positive scalar curvature.
+Let $(M^{4},g,f)$ be a non-compact, half-conformally flat (i.e., either $W^{+}=0$ or $W^{-}=0$), asymptotically conical Ricci expander with positive scalar curvature.
 By essentially the same arguments as in Chen-Wang [41] for gradient shrinking and steady Ricci solitons, one can show that $(M^{4},g,f)$ has vanishing $D$-tensor as introduced in [20, 21].
-Thus, by [19, Proposition 3.2], $(M^{4},g,f)$ is locally conformally flat.
+Thus, by [21, Proposition 3.2], $(M^{4},g,f)$ is locally conformally flat.
 Therefore, it follows from Theorem 3.1 that $(M^{4},g,f)$ has positive curvature operator $Rm>0$.
 ∎
-Finally, we conclude this section by proving Theorem 1.1.
+Finally, we conclude this section by providing a more direct proof of Theorem 1.1.
 Proof of Theorem 1.1.
-First, note that we can diagonalize the curvature operator $Rm$ with eigenvalues $m_{1}\leq m_{2}\leq m_{3}$. Then, the Ricci tensor $Rc$ is diagonalized with eigenvalues
+First, note that we can diagonalize the curvature operator $Rm$ with eigenvalues $m_{1}\leq m_{2}\leq m_{3}$, which are principal sectional curvatures of $(M^{3},g,f)$, with respect to a suitable orthonormal frame $\{e_{1},e_{2},e_{3}\}$ at each point $p\in M^{3}$.
+Then, the Ricci tensor $Rc$ is diagonalized with eigenvalues
 $(m_{1}+m_{2})\leq(m_{1}+m_{3})\leq(m_{2}+m_{3}),$
@@
 Assume instead that $m_{1}/R$ attains its negative infimum at infinity.
-By Lemma 6.1, any 3-dimensional cone with positive scalar curvature automatically has positive curvature operator/sectional curvature, and the smallest eigenvalue of its curvature operator is zero.
+By Lemma A.1, any 3-dimensional cone with positive scalar curvature automatically has positive curvature operator/sectional curvature, and the smallest eigenvalue of its curvature operator is zero.
 Then, as in Part I of the proof of Theorem 3.1, we get a contradiction.
@@
 Combining both cases, we conclude that $M^{-}=\emptyset$, so $M$ has nonnegative curvature operator or, equivalently, nonnegative sectional curvature (for $n=3$).
-Finally, by the same argument as in the proof of $Rm>0$ in Theorem 3.1, we conclude that $M^{3}$ has positive curvature operator. This completes the proof of Theorem 1.1. ∎
-Remark 3.2.
-Theorem 1.1 also follows from Theorem 1.2, which is valid for $n=3$ since the Weyl tensor vanishes automatically, and Lemma 6.1 on the curvature operator of 3D cones.
+Finally, by the same argument as in the proof of $Rm>0$ (i.e., Part III) for Theorem 3.1, we conclude that $M^{3}$ has positive curvature operator. This completes the proof of Theorem 1.1. ∎
 4. 4D asymptotically conical Ricci expanders with (half) PIC
@@
 By part (a), this implies $A_{1}\geq 0$.
-Following [24, Proposition 3.1(b)], we prove that $A_{1}>0$ by contradiction.
+Following [26, Proposition 3.1(b)], we prove that $A_{1}>0$ by contradiction.
 Suppose $A_{1}(p_{0})=0$ at some point $p_{0}\in M^{4}$.
@@
 Let $0\leq B_{1}\leq B_{2}\leq B_{3}$ be the singular eigenvalues of the matrix $B$ and ${\lambda}_{1}\leq{\lambda}_{2}\leq{\lambda}_{3}\leq{\lambda}_{4}$ be the eigenvalues of the Ricci tensor Rc{Rc}.
-Then, by [25, Lemma 2.2], the sum of the least two eigenvalues of $Rc$ is given by
+Then, by [27, Lemma 2.2], the sum of the least two eigenvalues of $Rc$ is given by
 $\lambda_{1}+\lambda_{2}=\tfrac{1}{2}(R-4B_{3}).$
 Thus, 2-nonnegative Ricci curvature is equivalent to $u:=R-4B_{3}\geq 0.$
-Now, by Lemma 2.1 and essentially the same computations as in the proof of [25, Theorem 3.1(a)], we have
+Now, by Lemma 2.1 and essentially the same computations as in the proof of [27, Theorem 3.1(a)], we have
 $\begin{split}\Delta_{f}u&=\Delta(R-4B_{3})\\ &\leq 2\rho u-\left[2|Rc|^{2}-8(A_{3}B_{3}+C_{3}B_{3}+2B_{1}B_{2})\right]\\ &\leq 2\rho u-8(A_{2}+A_{1}+C_{2}+C_{1})B_{3}.\end{split}$
 Then, by Lemma 2.1 and direct computations, for $F=f-2\log R$, we have
-$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&=\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{2}{R^{2}}\big[u|Rc|^{2}-4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\big],\end{split}$ (4.1)
+$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&=\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{2}{R^{2}}\big[u|Rc|^{2}-4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\big].\end{split}$ (4.1)
 Next, we prove, again by contradiction, that $(M^{4},g,f)$ has 2-nonnegative Ricci curvature. Suppose instead that the Ricci curvature is not 2-nonnegative. Then,
@@
 $u|Rc|^{2}-4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\equiv 0,$
-which implies $u|Rc|^{2}=4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}=0$ on $\Omega$.
+which implies $u|Rc|^{2}=4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\equiv 0$ on $\Omega$.
 Since $u(p_{0})<0$ by (4.2), we must have $|Rc|^{2}=0$ at $p_{0}$. But this is a contradiction to the scalar curvature $R>0$. Hence, Case 1 is ruled out.
@@
 Since $(M^{4},g,f)$ has PIC, the asymptotic cone must have WPIC.
-By Lemma 6.3, the Ricci curvature of the asymptotic cone is therefore nonnegative.
+By Lemma A.3, the Ricci curvature of the asymptotic cone is therefore nonnegative.
 Then by similar arguments in the proof of Theorem 3.1, we can rule out Case 2.
 Combining both cases, we conclude that $M^{-}=\emptyset$. Therefore, $M$ satisfies $u\geq 0$, completing the proof of 2-nonnegativity of the Ricci curvature.
-Finally, given that $(M^{4},g,f)$ has 2-nonnegative Ricci curvature, following the proof in [25, Theorem 3.1(b)], we shall prove 2-positive Ricci curvature by contradiction.
+Given that $(M^{4},g,f)$ has 2-nonnegative Ricci curvature, following the proof in [27, Theorem 3.1(b)], we shall prove 2-positive Ricci curvature by contradiction.
 We consider the quadratic form $Z:=RI-4\sqrt{B{}^{t}B}$, where $I$ is the 3 by 3 identity matrix.
@@
 where we have used the PIC condition and $B_{3}(p_{0})=\tfrac{R}{4}(p_{0})>0$ in the last inequality. Thus, we get a contradiction. Therefore, the Ricci curvature is 2-positive.
-This completes the proof of Theorem 1.4. ∎
-Finally, by Lemma 6.3 and an argument similar to that used in the proof of Theorem 1.4, we obtain the following slightly stronger result than Theorem 1.4.
+Finally, we establish the curvature estimate $|{Rm}|\leq 2R$.
+By the positive isotropic curvature condition, we have $|A|^{2}\leq\tfrac{3}{16}R^{2}$ and $|C|^{2}\leq\tfrac{3}{16}R^{2}$; see, e.g., the proof of [26, Theorem 1.3].
+Since the Ricci tensor of $(M^{4},g,f)$ is $2$-positive, it follows, as in the proof of [27, Theorem 3.1(a)], that $|{Rc}|^{2}\leq R^{2}$.
+On the other hand,
+$R^{2}\geq|{Rc}|^{2}=|\mathring{Rc}|^{2}+\tfrac{1}{4}R^{2}=4|B|^{2}+\tfrac{1}{4}R^{2}.$
+Therefore,
+$\begin{split}|{Rm}|^{2}&\leq 2\bigl(|A|^{2}+|C|^{2}+|B|^{2}\bigr)\leq\tfrac{9}{8}R^{2}.\end{split}$
+This completes the proof of the curvature bound, and hence the proof of Theorem 1.4. ∎
+Finally, by Lemma A.3 and an argument similar to that used in the proof of Theorem 1.4, we obtain the following slightly stronger result than Theorem 1.4.
 Theorem 4.2.
@@
   • (a)
-    If $(M^{4},g,f)$ satisfies $A_{2}\geq 0$ and $C_{2}\geq 0$, then the Ricci curvature is $2$-nonnegative.
+    If $(M^{4},g,f)$ satisfies $A_{2}\geq 0$ and $C_{2}\geq 0$, then the Ricci curvature is $2$-nonnegative, and $|Rm|\leq R$.
   • (b)
-    If $(M^{4},g,f)$ satisfies $A_{2}>0$ and $C_{2}>0$, then the Ricci curvature is $2$-positive.
+    If $(M^{4},g,f)$ satisfies $A_{2}>0$ and $C_{2}>0$, then the Ricci curvature is $2$-positive, and $|Rm|\leq R$.
 5. General lemma and further applications
-In this section, we formulate a fairly general method that can be effectively applied to prove generalized Hamilton-Ivey type curvature pinching estimates for a class of non-compact, asymptotically conical gradient Ricci solitons.
+In this section, we formulate a fairly general method that can be effectively applied to prove generalized Hamilton–Ivey type curvature pinching estimates for a class of non-compact, asymptotically conical gradient Ricci solitons.
 As applications, we obtain several analogues of other known curvature pinching results for ancient solutions, in the setting of asymptotically conical expanding Ricci solitons–including Theorem 1.5 as stated in the introduction, and Theorem 5.1 below.
@@
 Lemma 5.1.
-Let $(M^{n},g,f)$ be an $n$-dimensional non-compact gradient expanding Ricci soliton, satisfying Eq.
+Let $(M^{n},g,f)$ be an $n$-dimensional non-compact gradient Ricci soliton, satisfying Eq.
 (1.1) and with positive scalar curvature.
@@
 Let $F=f-2\log R$. Substituting (5.1) and the formula for $\Delta_{f}R$ in Lemma 2.1 into (5.3), we obtain
-$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&\leq\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{1}{R}(2\rho u-v)-\frac{u}{R^{2}}(2\rho R-2|Rc|^{2})\\ &\leq\frac{1}{R^{2}}(2u|Rc|^{2}-Rv).\end{split}$ (5.4)
+$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&=\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{1}{R}(2\rho u-v)-\frac{u}{R^{2}}(2\rho R-2|Rc|^{2})\\ &\leq\frac{1}{R^{2}}(2u|Rc|^{2}-Rv).\end{split}$ (5.4)
 Now, we prove by contradiction again. Suppose the lemma fails. Then the set
@@
 Remark 5.2.
-Lemma 5.1 also applies to gradient shrinking and steady Ricci solitons.
+Similarly, if a steady Ricci soliton $(M^{n},g,f)$ is asymptotic to a cylinder with $u\geq 0$, then $(M^{n},g,f)$ satisfies the asymptotic condition (5.2) in Lemma 5.1.
 5.2. The proof of Theorem 1.5
@@
 Proof of Theorem 1.5.
-First of all, it is easy to see that the inequality $B_{3}^{2}\leq A_{1}C_{1}$ implies nonnegative curvature operator $Rm\geq 0$ for $(M^{4},g)$; see, e.g., [43, Lemma 4.4] for a proof.
+First of all, it is easy to see that the inequality $B_{3}^{2}\leq A_{1}C_{1}$ implies nonnegative curvature operator $Rm\geq 0$ for $(M^{4},g)$; see, e.g., [44, Lemma 4.4] for a proof.
 Hence, to prove Theorem 1.5, it suffices to establish the inequality $B_{3}^{2}\leq A_{1}C_{1}$.
@@
 Step 1. To show $A_{3}\leq(6\Lambda^{2}+1)A_{1}$ and $C_{3}\leq(6\Lambda^{2}+1)C_{1}.$
-By the same computations as in [7, Lemma 6.1], we have
+By the same computations as in [8, Lemma 6.1], we have
 $\Delta_{f}[(6\Lambda^{2}+1)A_{1}-A_{3}]\leq 2\rho[(6\Lambda^{2}+1)A_{1}-A_{3}]-A_{3}^{2}.$
-Moreover, since the asymptotic cone $\mathcal{C}$ is a non-flat Euclidean cone, by (6.8) and (6.9), on the cone $\mathcal{C}$, we have
+Moreover, since the asymptotic cone $\mathcal{C}$ is a non-flat Euclidean cone, by (A.9) and (A.10), on the cone $\mathcal{C}$, we have
 $\bar{A}_{i}=\bar{C}_{j}=\bar{B}_{k},\quad 1\leq i,j,k\leq 3,$
@@
 $\gamma=\sup_{M}\frac{2B_{3}}{\sqrt{(A_{1}+A_{2})(C_{1}+C_{2})}}>1.$
-Let $w_{1}:=\tfrac{1}{2}\sqrt{(A_{1}+A_{2})(C_{1}+C_{2})}$. By the same computations as in [7, Lemma 6.2], we can find a positive constant $\delta_{1}>0$ such that
+Let $w_{1}:=\tfrac{1}{2}\sqrt{(A_{1}+A_{2})(C_{1}+C_{2})}$. By the same computations as in [8, Lemma 6.2], we can find a positive constant $\delta_{1}>0$ such that
 $\Delta_{f}(\gamma w_{1}-B_{3}-\delta_{1}R)\leq 2\rho(\gamma w_{1}-B_{3}-\delta_{1}R)-\delta_{1}|Rc|^{2}.$
@@
 $\gamma^{\prime}=\sup_{M}\frac{B_{3}}{\sqrt{A_{1}C_{1}}}>1.$
-Let $w_{2}:=\sqrt{A_{1}C_{1}}$. Then by the same computations as in [7, Proposition 6.3], we can find a positive constant $\delta_{2}>0$ such that
+Let $w_{2}:=\sqrt{A_{1}C_{1}}$. Then by the same computations as in [8, Proposition 6.3], we can find a positive constant $\delta_{2}>0$ such that
 $\Delta_{f}(\gamma^{\prime}w_{2}-B_{3}-\delta_{2}R)\leq 2\rho(\gamma^{\prime}w_{2}-B_{3}-\delta_{2}R)-\delta_{2}|Rc|^{2}.$
@@
 The assumption of the asymptotic cone being a (non-flat) Euclidean cone is used in the proofs of Step 2 and Step 3.
-Indeed, by (6.8) and (6.9), requiring either $4B_{3}^{2}\leq(A_{1}+A_{2})(C_{1}+C_{2})$ or $B_{3}^{2}\leq A_{1}C_{1}$ to hold on the asymptotic cone forces the link of the cone to be a (spherical) space form.
+Indeed, by (A.9) and (A.10), requiring either $4B_{3}^{2}\leq(A_{1}+A_{2})(C_{1}+C_{2})$ or $B_{3}^{2}\leq A_{1}C_{1}$ to hold on the asymptotic cone forces the link of the cone to be a (spherical) space form.
 5.3. Additional curvature pinching results
@@
   • (b)
-    Suppose $(M^{n},g,f)$ has WPIC1 and that the asymptotic cone has positive scalar curvature and WPIC2. Then, $(M^{n},g,f)$ must have WPIC2.
+    Suppose $(M^{n},g,f)$ has WPIC1555See, e.g., [3, Page 97] or [6, 11] for the definitions of WPIC1 and WPIC2.
+and that the asymptotic cone has positive scalar curvature and WPIC2.
+Then, $(M^{n},g,f)$ must have WPIC2.
   • (c)
@@
   • (d)
-    Suppose $(M^{n},g,f)$ is Kähler and has nonnegative orthogonal bisectional curvature. If the asymptotic cone has positive scalar curvature and WPIC2, then $(M^{n},g,f)$ must have WPIC2.
+    Suppose $(M^{n},g,f)$ is Kähler and has nonnegative orthogonal bisectional curvature [^6] See, e.g., [71, Page 28] for the definition of nonnegative orthogonal bisectional curvature..
+If the asymptotic cone has positive scalar curvature and WPIC2, then $(M^{n},g,f)$ must have WPIC2.
   • (e)
@@
 It suffices to verify, in each case, that the corresponding least curvature eigenvalue satisfies the differential equation (5.1) in Lemma 5.1.
-(a) By the same computations as in [72, Theorem 27] (see also [38, Lemma 2.4]), the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the Riemann curvature operator is satisfied.
+(a) By the same computations as in [73, Theorem 27] (see also [40, Lemma 2.4]), the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the Riemann curvature operator is satisfied.
 (b) By the same computations as in [3, Lemma 4.2], the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the complex sectional curvature is satisfied.
-(c) By the same computations as in [70, Proposition 4.2], the differential inequality (5.1) in Lemma 5.1 for the sum of the two least eigenvalues of the Ricci tensor is satisfied.
-(d) First of all, by the same computations as in [70, Lemma 6.1], the differential inequality (5.1) in Lemma 5.1 for the least bisectional curvature is satisfied.
+(c) By the same computations as in [71, Proposition 4.2], the differential inequality (5.1) in Lemma 5.1 for the sum of the two least eigenvalues of the Ricci tensor is satisfied.
+(d) First of all, by the same computations as in [71, Lemma 6.1], the differential inequality (5.1) in Lemma 5.1 for the least bisectional curvature is satisfied.
 Hence, by Lemma 5.1, the expanding Kähler-Ricci soliton has nonnegative bisectional curvature.
-Moreover, by the same arguments as in [71, Theorem 3.3], one can show that the smallest eigenvalue of the complex sectional curvature also satisfies the differential inequality (5.1) in Lemma 5.1.
-(e) Finally, by the same computations as in [43, Theorem 3.2] for $n\geq 12$ and in [40] for $9\leq n\leq 11$, the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the complex sectional curvature is satisfied.
+Moreover, by the same arguments as in [72, Theorem 3.3], one can show that the smallest eigenvalue of the complex sectional curvature also satisfies the differential inequality (5.1) in Lemma 5.1.
+(e) Finally, by the same computations as in [44, Theorem 3.2] for $n\geq 12$ and in [42] for $9\leq n\leq 11$, the differential inequality (5.1) in Lemma 5.1 for the least eigenvalue of the complex sectional curvature is satisfied.
 ∎
@@
 Other results that have been derived using B.-L.
-Chen’s lemma (see, e.g., [43, Corollary 2.4] or [25, Lemma 2.6]) for ancient solutions can similarly be extended to the setting of asymptotically conical gradient expanding Ricci solitons, following the same approach as above.
-6. Appendix
+Chen’s lemma (see, e.g., [44, Corollary 2.4] or [27, Lemma 2.6]) for ancient solutions can similarly be extended to the setting of asymptotically conical gradient expanding Ricci solitons, following the same approach as above.
+Appendix A Curvature properties of cones
 In this appendix, we examine the elementary relations between the curvature tensor of an $n$-dimensional cone, $n\geq 3$, with vanishing Weyl tensor and the curvature tensor of its link.
@@
 $Rm$ of $(\mathcal{C}^{n},g_{c})$ is given by
-$R_{1ijk}$ $=0,\quad 1\leq i,j,k\leq n,$ (6.1)
+$R_{1ijk}$ $=0,\quad 1\leq i,j,k\leq n,$ (A.1)
 $R_{abcd}$ $=r^{-2}\left[\bar{R}_{abcd}-\big(\bar{g}_{ac}\bar{g}_{bd}-\bar{g}_{ad}\bar{g}_{bc}\big)\right],\quad 2\leq a,b,c,d\leq n,$
 where $\bar{R}_{abcd}$ denotes the curvature tensor of the metric $\bar{g}$ of the link $\Sigma^{n-1}$. Moreover, the Ricci tensor $Rc$ of $(\mathcal{C}^{n},g)$ is given by
-$R_{1i}$ $=0,\quad 1\leq i\leq n,$ (6.2)
+$R_{1i}$ $=0,\quad 1\leq i\leq n,$ (A.2)
 $R_{ab}$ $=r^{-2}\left[\bar{R}_{ab}-(n-2)\bar{g}_{ab}\right],\quad 2\leq a,b\leq n.$
 and the scalar curvatures of $(\mathcal{C}^{n},g)$ and $(\Sigma^{n-1},\bar{g})$ are related by
-$R=r^{-2}\left[\bar{R}-(n-1)(n-2)\right].$ (6.3)
+$R=r^{-2}\left[\bar{R}-(n-1)(n-2)\right].$ (A.3)
 In addition, the nonzero Weyl curvature tensor $W$ of $(\mathcal{C}^{n},g)$ is given by
-$W_{1a1b}$ $=-\frac{1}{(n-2)r^{2}}\left(\bar{R}_{ab}-\frac{\bar{R}}{(n-1)}\bar{g}_{ab}\right),\quad 2\leq a,b\leq n,$ (6.4)
+$W_{1a1b}$ $=-\frac{1}{(n-2)r^{2}}\left(\bar{R}_{ab}-\frac{\bar{R}}{(n-1)}\bar{g}_{ab}\right),\quad 2\leq a,b\leq n,$ (A.4)
 $W_{abcd}$ $=r^{-2}\ \!\overline{W}_{abcd},\quad 2\leq a,b,c,d\leq n.$
-6.1. Curvature tensor of cones with vanishing Weyl tensor
+A.1. Curvature tensor of cones with vanishing Weyl tensor
 First, consider
@@
 In summary, we have the following basic fact in dimension $n=3$.
-Lemma 6.1.
+Lemma A.1.
 Let $\mathcal{C}^{3}:=[0,\infty)\times\Sigma^{2}$ be a $3$-dimensional cone equipped with the Riemannian metric $g_{c}=dr^{2}+r^{2}\bar{g}_{\Sigma}$.
@@
 where $(\Sigma^{n-1},\bar{g})$ is a closed (n1)(n-1)-dimensional Riemannian manifold.
-Then, by (6.4), we see that $\mathcal{C}$ has vanishing Weyl curvature $W=0$ if and only if its link $\Sigma$ is a space form.
+Then, by (A.4), we see that $\mathcal{C}$ has vanishing Weyl curvature $W=0$ if and only if its link $\Sigma$ is a space form.
 Thus, we immediately have the following
-Lemma 6.2.
+Lemma A.2.
 Let $\mathcal{C}^{n}:=[0,\infty)\times\Sigma^{n-1}$ be an $n$-dimensional ($n\geq 4$) cone with nonnegative scalar curvature, equipped with the Riemannian metric $g_{c}=dr^{2}+r^{2}\bar{g}_{\Sigma}$.
@@
 In particular, if $(\operatorname{\mathcal{C}}^{n},g_{c})$ is locally conformally flat and has nonnegative scalar curvature, then it has nonnegative curvature operator $Rm\geq 0$.
-6.2. Curvature decomposition and curvature operator of 4D cones
+A.2. Curvature decomposition and curvature operator of 4D cones
 Recall that, with respect to the decomposition
@@
 on any oriented smooth Riemannian 4-manifold $(M^{4},g)$, the curvature operator of $(M^{4},g)$ admits the following decomposition:
-$Rm=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\frac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\frac{R}{12}I\end{pmatrix}$, (6.5)
+$Rm=\begin{pmatrix}A&B\\ B^{t}&C\end{pmatrix}=\begin{pmatrix}W^{+}+\frac{R}{12}I&\mathring{Rc}\\ \mathring{Rc}&W^{-}+\frac{R}{12}I\end{pmatrix}$, (A.5)
 where $W^{\pm}$ denote the self-dual and anti-self-dual Weyl curvature tensors, respectively, and $\mathring{Rc}$ denotes the traceless Ricci tensor.
@@
 The inner product on 2-forms is defined by
-$\langle X\wedge Y,V\wedge W\rangle=\langle X,V\rangle\langle Y,W\rangle-\langle X,W\rangle\langle Y,V\rangle.$
-Observe that, for the matrices $A$ and $C$ in (6.5), we have
+$\langle X\wedge Y,V\wedge W\rangle=\langle X,V\rangle\langle Y,W\rangle-\langle X,W\rangle\langle Y,V\rangle.$ (A.6)
+Observe that, for the matrices $A$ and $C$ in (A.5), we have
 […]
@@
 of $\wedge^{2}T\Sigma^{3}$, where $\{\bar{e}_{2},\bar{e}_{3},\bar{e}_{4}\}$ is a local orthonormal frame of $T\Sigma^{3}$. Suppose that, in this frame, $\overline{Rm}$ is diagonal with entries
-$\bar{R}_{2323}=:m_{1},\quad\bar{R}_{2424}=:m_{2},\quad\bar{R}_{3434}=:m_{3}$
+$\bar{R}_{2323}=:m_{3},\quad\bar{R}_{2424}=:m_{2},\quad\bar{R}_{3434}=:m_{1}$
 such that $m_{1}\leq m_{2}\leq m_{3}$. Then, with respect to the tangent frame $\{\bar{e}_{2},\bar{e}_{3},\bar{e}_{4}\}$, the Ricci tensor $\overline{Rc}$ of $\Sigma^{3}$ takes the form
@@
 $e_{1}=\partial_{r},\qquad e_{i}=r^{-1}\bar{e}_{i}\ (i=2,3,4),$
-so that $\{e_{1},e_{2},e_{3},e_{4}\}$ forms an orthonormal frame of $T\mathcal{C}$ with respect to $g_{c}$. Then, by (6.1), we have
+so that $\{e_{1},e_{2},e_{3},e_{4}\}$ forms an orthonormal frame of $T\mathcal{C}$ with respect to $g_{c}$. Then, by (A.1), we have
 $R_{1jkl}$ $=0,\quad 1\leq j,k,l\leq 4,$
@@
 where $\bar{R}_{abcd}$ denotes the curvature tensor of the link $(\Sigma^{3},\bar{g})$.
-Therefore, by (6.6), we obtain
-$A_{11}=C_{11}$ $=\tfrac{1}{2}R_{3434}=\tfrac{1}{2r^{2}}(m_{3}-1),$ (6.8)
+Therefore, by (A.7), we obtain
+$A_{11}=C_{11}$ $=\tfrac{1}{2}R_{3434}=\tfrac{1}{2r^{2}}(m_{1}-1),$ (A.9)
 $A_{22}=C_{22}$ $=\tfrac{1}{2}R_{2424}=\tfrac{1}{2r^{2}}(m_{2}-1),$
-$A_{33}=C_{33}$ $=\tfrac{1}{2}R_{2323}=\tfrac{1}{2r^{2}}(m_{1}-1),$
+$A_{33}=C_{33}$ $=\tfrac{1}{2}R_{2323}=\tfrac{1}{2r^{2}}(m_{3}-1),$
 and $A_{ij}=C_{ij}=0\ (i\neq j)$, e.g.,
 $A_{12}=C_{12}=\tfrac{1}{2}R_{3442}=0,\quad\text{etc.}$
-Similarly, by (6.7), we have
-$B_{11}$ $=-\tfrac{1}{2}R_{3434}=-\tfrac{1}{2r^{2}}(m_{3}-1)=-A_{11},$ (6.9)
+Similarly, by (A.8), we have
+$B_{11}$ $=-\tfrac{1}{2}R_{3434}=-\tfrac{1}{2r^{2}}(m_{1}-1)=-A_{11},$ (A.10)
 $B_{22}$ $=-\tfrac{1}{2}R_{4242}=-\tfrac{1}{2r^{2}}(m_{2}-1)=-A_{22},$
-$B_{33}$ $=-\tfrac{1}{2}R_{2323}=-\tfrac{1}{2r^{2}}(m_{1}-1)=-A_{33},$
+$B_{33}$ $=-\tfrac{1}{2}R_{2323}=-\tfrac{1}{2r^{2}}(m_{3}-1)=-A_{33},$
 and $B_{ij}=0\ (i\neq j)$, e.g.,
@@
 Moreover, for the Ricci tensor of $(\mathcal{C},g_{c})$, we have $R_{ij}=0\ (i\neq j)$ and
-$R_{11}$ $=0,$ $R_{22}=r^{-2}\left(m_{1}+m_{2}-2\right),$
-$R_{33}$ $=r^{-2}\left(m_{1}+m_{3}-2\right),$ $R_{44}=r^{-2}\left(m_{2}+m_{3}-2\right).$
+$R_{11}$ $=0,$ $R_{22}=r^{-2}\left(m_{2}+m_{3}-2\right),$
+$R_{33}$ $=r^{-2}\left(m_{1}+m_{3}-2\right),$ $R_{44}=r^{-2}\left(m_{1}+m_{2}-2\right).$
 In conclusion, based on the above computations, we have
-Lemma 6.3.
+Lemma A.3.
 Let $\mathcal{C}^{4}:=[0,\infty)\times\Sigma^{3}$ be a $4$-dimensional cone equipped with the Riemannian metric $g_{c}=dr^{2}+r^{2}\bar{g}_{\Sigma}$. Then, the following statements hold:
@@
   • (ii)
-    $(\mathcal{C}^{4},g_{c})$ has half WPIC if and only if it has WPIC and nonnegative Ricci curvature $Rc\geq 0$, if and only if $\overline{Rc}\geq 2{\bar{g}}$.
+    $(\mathcal{C}^{4},g_{c})$ has half-WPIC if and only if it has WPIC and nonnegative Ricci curvature $Rc\geq 0$, if and only if $\overline{Rc}\geq 2{\bar{g}}$.